### v1

# Matching conditions at null infinity in the presence of logarithms: the role of advanced and retarded radiation

Abstract
We provide a new perspective on the general matching conditions between the future of past null infinity and the past of future null infinity, emphasizing the impact of dominant logarithmic terms in the asymptotic expansion of the fields near null infinity. We explicitly consider the cases of a massless scalar field and of electromagnetism. Key in our derivation is the identification of the physical origin of these logarithms, which are associated with advanced and retarded radiation saturating the finite energy flux condition at null infinity (in a space of functions which is made precise). The matching conditions arise then from the requirement of Coulombic (i.e., $1/r$) behaviour at spatial infinity.
1 Introduction
Matching conditions between the future of past null infinity and the past of future null infinity play a central role in the study of conservation laws and scattering processes involving massless fields Strominger:2013jfa ; He:2014cra ; Kapec:2015ena ; Strominger:2017zoo . As explained in Strominger:2017zoo (in particular text above (5.2.10)), these matching conditions rely on a certain number of assumptions concerning the decay of the fields at null infinity and their behaviour near its boundaries. This paper investigates from a physical viewpoint the matching conditions in a more general context when these assumptions are relaxed, both for a massless scalar field and for free electromagnetism in four dimensional Minkowski space.
To formulate the problem in a concrete manner, let us consider the definite case of a massless scalar field, denoted by $\phi$. The standard behaviour assumed near future null infinity is $\phi\sim\frac{a(u,x^{A})}{r}$ where $a(u,x^{A})=O(1)$ as $u\rightarrow$-$\infty$, i.e., $\lim_{u\rightarrow-\infty}a(u,x^{A})=A(x^{A})$ for some definite function $A(x^{A})$ on the sphere. Here, $u$ is the retarded time trt-r and $x^{A}$ are angular coordinates. Similarly, one assumes near past null infinity (with $v$ the advanced time $t+r$) that $\phi\sim\frac{b(v,x^{A})}{r}$ with $\lim_{v\rightarrow\infty}b(v,x^{A})=B(x^{A})$ for some definite function $B(x^{A})$ on the sphere. The matching conditions for the scalar field are $A(x^{A})=B(-x^{A})$ where $x^{A}\rightarrow-x^{A}$ stands for the antipodal map Campiglia:2017dpg ; Campiglia:2017xkp . The matching conditions for electromagnetism and gravity take similar forms. Further matching conditions involving the subleading orders and associated with an infinite tower of charges have been investigated in Briceno:2025ivl .
A spatial-infinity-based derivation of the matching conditions was given in Henneaux:2018gfi ; Henneaux:2018hdj ; Henneaux:2018mgn , where they were shown to follow from (and in fact be equivalent with) the assumption that the leading orders of the field in the asymptotic expansion near spatial infinity should obey appropriate parity conditions under the sphere antipodal map. These parity conditions make the action finite “on the nose” (without need for regularization). In the scalar field case, the parity conditions read explicitly, in terms of the field $\phi(r,x^{A})$ and its conjugate $\pi(r,x^{A})$ on the initial slice t=0t=0111These conditions are invariant under Poincaré transformations.,
$\phi=\frac{C(x^{A})}{r}+o(r^{-1})\,,\qquad\pi=\frac{P(x^{A})}{r^{2}}+o(r^{-2})$ (1)
with
$C(-x^{A})=C(x^{A})\,,\qquad P(-x^{A})=-P(x^{A})\,.$ (2)
The reason that these parity conditions on the initial data are equivalent to the matching conditions of Strominger:2013jfa ; He:2014cra ; Kapec:2015ena ; Strominger:2017zoo can be traced to an interesting property of the asymptotic hyperbolic coordinates that connect spacelike infinity to null infinity Ashtekar:1978zz ; Beig:1982ifu , namely, that parity properties under the sphere antipodal map on Cauchy hyperplanes become parity properties under the hyperboloid antipodal map, involving not only the sphere antipodal map but also an hyperbolic time inversion that reverses past and future Compere:2011ve ; Troessaert:2017jcm .
Now, although very general Christodoulou:1993uv , the behaviour assumed in Strominger:2013jfa ; He:2014cra ; Kapec:2015ena ; Strominger:2017zoo near the past (future) of future (past) null infinity is not the most general one that is compatible with the condition of finite energy flux through null infinity. Taking again the example of a free massless scalar field in Minkowski space, one can show that the total energy radiated through $\mathscr{I}^{+}$ is given by Frolov:1977bp
$\left.\Delta E\right|_{\mathscr{I}^{+}}=-\int_{\mathscr{I}^{+}}dud^{2}\hat{x}\sqrt{\gamma}\left(\partial_{u}a\right)^{2}\,.$ (3)
Here $\gamma$ denotes the determinant of the metric of the round 2-sphere $\gamma_{AB}$. This allows a logarithmic behaviour of $a(u)$ in the limit $u\rightarrow$-$\infty$, which is clearly more singular than the $O(1)$-behaviour mentioned above [^2] If we assume that in the limit $u\rightarrow$-$\infty$, $\partial_{u}a=O(u^{k})$ where $k\in\mathbb{Z}$ is an integer, then convergence holds for $k\leq-1$. We shall say that the upper value k=1k=-1 saturates the finite energy flux condition. It is natural to adopt $k\in\mathbb{Z}$ because fractional powers of $u$ would yield fractional powers of $r$ at spatial infinity, which we exclude in this paper..
The question is then: what do the matching conditions become if one includes this more general logarithmic behaviour? The answer to that question was given in Fuentealba:2024lll ; Fuentealba:2025ekj by following again a spatial-infinity-based route in which one integrates the field equations from initial data that keep the same standard $1/r$ behaviour of the elementary solution of the Poisson equation at spatial infinity, but with leading coefficients that do not obey definite parity condition and contain instead both parities (i.e., no parity restriction). More explicitly, in the scalar field case, one keeps (1) – which we call “Coulomb behaviour” even when the coefficients $C$ and $P$ depend on the angles –, but one drops (2). Such initial data were shown to lead to the above $\log(-u)/r$ behaviour near the past of future null infinity, which is paired with a leading $\log r/r$ near null infinity. The coefficients of these $\log(-u)/r$ and $\log r/r$ logarithmic terms come from the other parity component at spatial infinity and obey therefore opposite matching conditions to those of Strominger:2013jfa ; He:2014cra ; Kapec:2015ena ; Strominger:2017zoo . Such mixed matching conditions were found earlier in higher dimensions in Henneaux:2019yqq .
The purpose of this paper is to provide new insight on the matching conditions through a different, physically motivated explanation for the emergence of logarithmic terms in the asymptotic expansion at null infinity of massless scalar and electromagnetic fields in Minkowski space. Our approach is complementary to that of Fuentealba:2024lll ; Fuentealba:2025ekj in that it relies from the very begining on the behaviour of the fields near null infinity rather than on their Cauchy development from initial data.
We show that logarithmic terms at future null infinity originate from advanced waves that saturate the physically sensible condition that the total energy entering the spacetime through past null infinity is finite. This requirement yields an advanced solution exhibiting a $\log\left(v\right)$ behavior at late advanced times (rather than the stronger $O(1)$ behaviour), which, in turn, leaves a distinct imprint at future null infinity characterized by a $\log\left(r\right)/r$ decay near $\mathscr{I}^{+}$, which dominates the usually assumed $r$-asymptotic expansion. Analogously, logarithmic terms at past null infinity originate from retarded waves exhibiting a $\log\left(-u\right)$ behavior at early retarded times (see Figure 1). The asymptotic conditions on the fields at spatial infinity, which are requested to keep the Coulombic behaviour (1), connects the advanced and retarded branches, from which one reads the matching conditions. These coincide with those of Fuentealba:2024lll ; Fuentealba:2025ekj , of which it provides an alternative derivation.
Figure 1: The double cover of the Penrose diagram for Minkowski space depicting, in blue, the trace of the last advanced modes at future null infinity $\mathscr{I}^{+}$ and, in red, the first retarded modes at past null infinity $\mathscr{I}^{-}$. The advanced branch corresponds to solutions whose logarithmic falloff $\log(r)/r$ at $\mathscr{I}^{+}$ originates from finite-energy data entering through $\mathscr{I}^{-}$, while the retarded branch encodes the matching conditions at early times. Together, they illustrate how logarithmic terms at null infinity naturally emerge from the interplay between advanced and retarded solutions saturating the finite-energy condition.
The connection between advanced waves and non-power-law radial expansions of the fields near $\mathscr{I}^{+}$ violating the peeling property of $\mathscr{I}^{+}$ has been recognized since the earliest studies of gravitational radiation at null infinity. In fact, the standard power-law expansion of the leading orders in the asymptotic expansion in the BMS formalism might be considered to be equivalent to Sommerfeld’s radiation condition, which is required to describe radiation emitted by bounded sources (see, e.g., Bondi:1962px ; Sachs:1961zz ; Sachs:1962wk ; Tamburino:1966zz ; Kroon:1998dv ). For this reason, such terms were typically neglected due to the imposition of the Sommerfeld radiation condition. Nonetheless, there could be physically relevant scenarios where incoming radiation, and consequently leading logarithmic terms, could play a crucial role. This occurs for instance in the study of scattering processes Kraus:2025wgi .
We go beyond these earlier observations on the connection between non-polynomial terms at null infinity and advanced radiation by making the explicit connection with the initial data on Cauchy hyperplanes and the resulting Cauchy development, obtaining thereby a new physical insight on the generalized matching conditions valid when these logarithmic terms, involving both $\log(-u)$ and $\log r$, are present.
Our paper is organized as follows. In the next section (Section 2), we treat the case of the scalar field. We cover in detail monopole and dipole radiation fields, which exhibit the main points. We then move in Section 3 to the electromagnetic case, of which we directly provide the all-multipole analysis. We first give the form of the electromagnetic potential and of the electromagnetic field. We investigate next the conditions under which the energy fluxes through past and future infinity are finite. These conditions, together with the requested power law behaviour at spatial infinity, lead then to the matching conditions for electromagnetism. Section 4 is devoted to conclusions and prospects. Three appendices complete our paper: Appendix A covers the case of a scalar field with a general multipole expansion. Appendix B provides a thorough derivation of the general solution of Maxwell’s equations in retarded coordinates. Finally, Appendix C lists useful identities on infinite sums associated with the Legendre polynomials.
2 Massless scalar field
We start with a massless scalar field in four spacetime dimensions, obeying the wave equation
$\Box\phi=0\,$, (4)
which reads explicitly, in retarded coordinates $(u=t-r,r,x^{A})$,
$2\partial_{u}\partial_{r}\phi+\partial_{r}^{2}\phi+\frac{2\partial_{r}\phi}{r}-\frac{2\partial_{u}\phi}{r}+\frac{D^{2}\phi}{r^{2}}=0\,$ (5)
where $D^{2}$ is the Laplacian on the unit $2$-sphere.
We impose the following asymptotic conditions:
  • The field can be written as the sum of retarded and advanced waves
    $\phi=\frac{\bar{\phi}_{R}\left(u,x^{A}\right)}{r}+\frac{\bar{\phi}_{A}\left(v,x^{A}\right)}{r}+O(r^{-2})\,$, (6)
    where the functions $\bar{\phi}_{R}$ and $\bar{\phi}_{A}$, which are smooth but not necessarily analytic in u=tru=t-r or $v=t+r$, are required to give finite energy fluxes through null infinity. The fact that we write the field $\phi$ as a sum of a retarded part and an advanced one does not mean, of course, that $u$ and $v$ are independent variables (given $r$), but simply that there is one part of $\phi$ that is a function of time through $u$ and the other, through $v$. When computing the wave equation in retarded coordinates (say), one should of course replace $v$ by $u+2r$ (and vice-versa, use u=v2ru=v-2r when working in advanced coordinates).
  • The field $\phi$ has the “Coulombic” $1/r$ behaviour at spatial infinity, characteristic of the elementary solution of Poisson’s equation, i.e., on constant Minkowskian time hyperplanes,
    $\phi=\frac{C(x^{A})}{r}+o(r^{-1})\,$, (7)
    and its conjugate momentum behaves as
    $\pi=\frac{P(x^{A})}{r^{2}}+o(r^{-2})\,$, (8)
    in agreement with (1). But note that coefficients that depend on the angles are allowed.
2.1 Spherical waves
2.1.1 Logarithmic terms at null infinity from advanced and retarded radiation
We first illustrate the approach by considering the simplest case of a spherical wave. In this case, the functions $\bar{\phi}_{R}$ and $\bar{\phi}_{A}$ do not depend on the angles and the general solution to the wave equation reduces to [^3] Note that the split of the static component $\phi=\frac{a}{r}$ is ambiguous since it corresponds to $\bar{\phi}_{R}\left(u\right)=a-\lambda$, $\bar{\phi}_{A}\left(v\right)=\lambda$, with $\lambda$ arbitrary. We will lift this ambiguity by imposing that the constant $a$ be equally split between the retarded and advanced parts, i.e., $\lambda=\frac{a}{2}$.
$\phi=\frac{\bar{\phi}_{R}\left(u\right)}{r}+\frac{\bar{\phi}_{A}\left(v\right)}{r}\,.$ (9)
It is important to emphasize that the functions $\bar{\phi}_{R}\left(u\right)$ and $\bar{\phi}_{A}\left(v\right)$ are not determined by the differential equation and are completely arbitrary. Nonetheless, this freedom must be partially constrained by the fundamental physical requirement that the total radiated energy be finite.
As indicated above, an expression for the total radiated or absorbed energy was derived in Ref. Frolov:1977bp , by projecting the energy-momentum tensor of the scalar field onto $\mathscr{I}^{\pm}$. In particular, for the spherically symmetric solution given in Eq. (9), the total energy radiated through $\mathscr{I}^{+}$ by the retarded field is given by
$\left.\Delta E\right|_{\mathscr{I}^{+}}=-\int_{\mathscr{I}^{+}}dud^{2}\hat{x}\sqrt{\gamma}\left(\partial_{u}\bar{\phi}_{R}\right)^{2}\,.$ (10)
To ensure that the total energy radiated by the retarded field remains finite, in the limit $u\rightarrow\pm\infty$, the behavior of the field $\bar{\phi}_{R}$ near $\mathscr{I}_{+}^{+}$ and $\mathscr{I}_{-}^{+}$ must be, at most, of the form (in the space of functions such that $\partial_{u}f\sim u^{k}$, $k\in\mathbb{Z}$):
$\bar{\phi}_{R}\left(u\right)\underset{u\rightarrow\pm\infty}{=}\phi_{R}^{\pm}\log\left(\pm u\right)+\varphi_{R}^{\pm}+o(1),$ (11)
where the coefficients $\phi_{R}^{\pm}$ and $\varphi_{R}^{\pm}$ are constants for the spherical wave.
Analogously, the total radiation entering the spacetime through $\mathscr{I}^{-}$ is
$\left.\Delta E\right|_{\mathscr{I}^{-}}=\int_{\mathscr{I}^{-}}dvd^{2}\hat{x}\sqrt{\gamma}\left(\partial_{v}\bar{\phi}_{A}\right)^{2}\,.$ (12)
Therefore, the condition for having a finite total incoming energy is
$\bar{\phi}_{A}\left(v\right)\underset{v\rightarrow\pm\infty}{=}\phi_{A}^{\pm}\log\left(\pm v\right)+\varphi_{A}^{\pm}+o(1)\,$, (13)
where $\phi_{A}^{\pm}$ and $\varphi_{A}^{\pm}$ are constants for the spherical wave.
The key step is to rewrite Eq. (9) entirely in terms of the retarded time, $u$ using the relation $v=u+2r$. In this form, the general solution for a spherically symmetric wave becomes:
$\phi=\frac{\bar{\phi}_{R}\left(u\right)}{r}+\frac{\bar{\phi}_{A}\left(u+2r\right)}{r}.$ (14)
Taking the limit to future null infinity ($u=\text{const}$ and $r\rightarrow\infty$), the argument of the advanced component of the solution is dominated by $r$. Thus, using the asymptotic behavior (13), one finds
$\left.\phi\right|_{\mathscr{I}^{+}}=\phi_{A}^{+}\frac{\log r}{r}+\left(\bar{\phi}_{R}\left(u\right)+\phi{}_{A}^{+}\log\left(2\right)+\varphi_{A}^{+}\right)\frac{1}{r}+o(r^{-1})\,$, (15)
which is precisely the logarithmic behavior described in Fuentealba:2024lll
$\left.\phi\right|_{\mathscr{I}^{+}}=\frac{1}{2}\Psi\frac{\log\left(r\right)}{r}+\frac{\Phi}{r}+o(r^{-1})\,$, (16)
with
$\Phi=\bar{\phi}_{R}\left(u\right)+\phi{}_{A}^{+}\log\left(2\right)+\varphi_{A}^{+}\,$, (17)
and
$\Psi=2\phi_{A}^{+}.$
This analysis shows that the late-time behavior of the advanced wave manifests itself at future null infinity as a term of the form $\log\left(r\right)/r$. One might fear that this $\frac{\log\left(r\right)}{r}$-term will jeopardize the finite energy flux condition due to its slowlier decay at null infinity, but this is not the case because it does not depend on $u$ so that its $u$-derivative vanishes. Note also that the term of order $r^{-1}$ acquires a time-independent component from the advanced solution, which, due to its independence of $u$ , is not part of the radiative field at future null infinity. The other $\frac{1}{r}$-contributions from the advanced wave depend on $u$, but are subleading in the limit $u\rightarrow$-$\infty$ (decay as $\sim u^{-1}$ or faster Fuentealba:2024lll ).
Analogously, if we perform a similar analysis at past null infinity $\mathscr{I}^{-}$ one finds
$\left.\phi\right|_{\mathscr{I}^{-}}=\phi_{R}^{-}\frac{\log r}{r}+\left(\bar{\phi}_{A}\left(v\right)+\phi_{R}^{-}\log\left(2\right)+\varphi_{R}^{-}\right)\frac{1}{r}+o(r^{-1})\,.$
Therefore, in the expansion at past null infinity corresponding to (16)
$\left.\phi\right|_{\mathscr{I}^{-}}=\frac{1}{2}\Psi^{\prime}\frac{\log\left(r\right)}{r}+\frac{\Phi^{\prime}}{r}+o(r^{-1})\,$, (18)
one identifies
$\Phi^{\prime}$ $=\bar{\phi}_{A}\left(v\right)+\phi_{R}^{-}\log\left(2\right)+\varphi_{R}^{-}\,$, (19)
and
$\Psi^{\prime}=2\phi_{R}^{-}.$
2.1.2 Matching conditions
We now derive the matching conditions for spherical waves near spatial infinity $i^{0}$.
For the retarded solution, spatial infinity is approached in the limit $u\rightarrow$-$\infty$ with large $r$, such that the combination $t=u+r$ remains finite. Using Eq. (11) one finds
$\lim_{u\rightarrow-\infty}\frac{\bar{\phi}_{R}}{r}$ $=\lim_{r\rightarrow+\infty}\frac{\phi_{R}^{-}\log\left(r-t\right)}{r}+\frac{\varphi_{R}^{-}}{r}=\phi_{R}^{-}\frac{\log\left(r\right)}{r}+\frac{\varphi_{R}^{-}}{r}+\dots\,$,
Analogously, the limit of the advanced solution to spatial infinity is given by the limit $v\rightarrow+\infty$ with large $r$, such that the combination t=vrt=v-r remains finite. Therefore, using Eq. (13) one finds
$\lim_{v\rightarrow+\infty}\frac{\bar{\phi}_{A}}{r}$ $=\lim_{v\rightarrow+\infty}\frac{\phi_{A}^{+}\log\left(r+t\right)}{r}+\frac{\varphi_{A}^{+}}{r}=\frac{\phi_{A}^{+}\log\left(r\right)}{r}+\frac{\varphi_{A}^{+}}{r}+\dots\,.$
To obtain the full solution at $i^{0}$, we must consider the sum of the contributions from the advanced and retarded waves. Consequently, the leading order of the solution near $i^{0}$ is given by
$\left.\phi\right|_{i^{0}}=\left(\phi_{R}^{-}+\phi_{A}^{+}\right)\frac{\log\left(r\right)}{r}+\frac{\varphi_{A}^{+}+\varphi_{R}^{-}}{r}+\dots\,.$
As recalled in the introduction and argued by many authors, the natural behaviour at spatial infinity is the one of the elementary solution of Poisson equation, i.e., $\phi\sim\frac{1}{r}$. Thus, the $\log r$ terms should cancel in the expansion of the field near $i^{0}$. This implies that one must impose the following condition:
$\phi_{R}^{-}=-\phi_{A}^{+}.$ (20)
In terms of the variables defined in Eqs. (16) and (18), this becomes
$\Psi=-\Psi^{\prime},$
which precisely coincides with the matching condition established in Fuentealba:2024lll for the coefficient of the leading logarithmic $\log r/r$ term, applied to the spherical wave. This matching condition corresponds to the $Q$ branch in the language of Fuentealba:2024lll and fulfills matching conditions with the minus sign, opposite to the matching condition adopted in Strominger:2013jfa ; He:2014cra ; Kapec:2015ena ; Strominger:2017zoo and valid for the subleading $P$-branch.
It is through the matching conditions that the coefficients of the $\frac{\log r}{r}$-term and of the $\frac{\log u}{r}$-term in the expansion near future null infinity are forced to be equal (up to the sign). They would otherwise be unrelated, having a priori different origins.
The $P$-branch is the dominant one if one assumes from the outset that there is no leading $\log r/r$ term in the expansion of the scalar field near future null infinity, i.e., $\phi_{A}^{+}=0$ (and thus also $\phi_{R}^{-}=0$ by (20)). In that case one finds that $\bar{\phi}_{R}\left(u\right)$ (respectively, $\bar{\phi}_{A}\left(v\right)$) behaves near the past of future null infinity (respectively, the future of past null infinity) as
$\bar{\phi}_{R}\left(u\right)=\varphi_{R}^{-}\,,\qquad\bar{\phi}_{A}\left(v\right)=\varphi_{A}^{+}$ (21)
up to terms that vanish in the limit. Hence, one has
$\Psi=\Psi^{\prime}=0\,,\qquad\lim_{u\rightarrow-\infty}\Phi=\varphi_{R}^{-}+\varphi_{A}^{+}=+\lim_{v\rightarrow\infty}\Phi^{\prime}\,$, (22)
(and $\left.\phi\right|_{i^{0}}=\frac{\varphi_{A}^{+}+\varphi_{R}^{-}}{r}+\dots$) which is the matching condition, with a positive sign, for the $P$-branch Campiglia:2017dpg . It is of interest to emphasize that the matching conditions, which involve the sum $\varphi_{A}^{+}+\varphi_{R}^{-}$, are insensitive to the ambiguity in the decomposition of the static Coulomb part between retarded and advanced solutions.
As explained in Fuentealba:2024lll , when the $Q$-branch is present ($\Psi\not=0$, $\Psi^{\prime}\not=0$), one must substract its contributions to the $1/r$-term in order to dig out the $P$-contributions, which obey (22).
We stress that the retarded and advanced solutions, when expressed in terms of the retarded or advanced coordinate, have a purely power-law behaviour in the radial coordinate (in this case just $1/r$), with no logarithmic term involving $r$. The logarithms at future null infinity appear only when the advanced solution is expressed in terms of the retarded time (and similarly for past null infinity). For that reason, the logarithms at future null infinity can be eliminated by assuming no advanced radiation. By the matching at spatial infinity, this would also eliminate the $\log(-u)$ term in $\bar{\phi}_{R}$.
2.2 Dipole waves
2.2.1 Logarithmic terms at null infinity from advanced and retarded radiation
Since the equations become increasingly involved when one includes higher spherical harmonics, we shall treat explicitly the dipole case in the core of the text. It provides a very good illustration of the general case, which is covered in Appendix A.
The general solution of the wave equation in a source-free region can again be expressed as the sum of retarded and advanced terms.
$\phi=\phi_{R}+\phi_{A},$
where we have now, following Bondi:1962px ,
$\phi_{R}$ $=\sum_{m=-1,0,1}\left(\frac{1}{r}\frac{da_{m}\left(u\right)}{du}-\frac{1}{r^{2}}a_{m}\left(u\right)\right)Y_{1m}\left(\theta,\phi\right),$ (23)
$\phi_{A}$ $=\sum_{m=-1,0,1}\left(\frac{1}{r}\frac{db_{m}\left(v\right)}{dv}-\frac{1}{r^{2}}b_{m}\left(v\right)\right)Y_{1m}\left(\theta,\phi\right).$ (24)
Here, $a_{m}\left(u\right)$ and $b_{m}\left(v\right)$ are arbitrary functions of the retarded and advanced times, respectively. We find again the same ambiguity in the decomposition between retarded and advanced parts of the Coulomb static component, now at dipole order $1/r^{2}$.
[…]
$\phi_{R}=\frac{\bar{\phi}_{R}\left(u,\hat{x}\right)}{r}+O\left(r^{-2}\right),$ (25)
with
$\bar{\phi}_{R}\left(u,\hat{x}\right)=\sum_{m}\frac{da_{m}\left(u\right)}{du}\,Y_{1m}\,.$ (26)
Finiteness of the total energy flux implies therefore the condition
$\lim_{u\rightarrow\pm\infty}\frac{da_{m}\left(u\right)}{du}=\phi_{R\left(m\right)}^{\pm}\log\left(\pm u\right)+\varphi_{R\left(m\right)}^{\pm}+\dots\,$, (27)
which leads upon integration, near the past of future null infinity ($u\rightarrow$-$\infty$) to
$a_{m}\left(u\right)=\phi_{R\left(m\right)}^{-}\left(u\log(-u)-u\right)+\varphi_{R\left(m\right)}^{-}u+c_{R\left(m\right)}+\dots$ (28)
where the dots denote subleading terms (in $u$). Similarly, the condition of finiteness of the total flux through past infinity implies ($v\rightarrow\infty$)
$b_{m}\left(v\right)=\phi_{A\left(m\right)}^{+}\left(v\log(v)-v\right)+\varphi_{A\left(m\right)}^{+}v+c_{A\left(m\right)}+\dots$ (29)
Substituting Eq. (29) into Eq. (24), replacing $v$ by $u+2r$ and taking the limit $r\rightarrow\infty$ with $u$ held constant (which implies $v\rightarrow\infty$), one finds that the solution can then be written near $\mathscr{I}^{+}$ as:
$\left.\phi\right|_{\mathscr{I}^{+}}$ $=-\left[\sum_{m}\phi_{A\left(m\right)}^{+}Y_{lm}\right]\frac{\log r}{r}$
$+\left[\sum_{m}\left(\frac{d}{du}a_{m}\left(u\right)-\left(\phi_{A\left(m\right)}^{+}\log 2+\varphi_{A\left(m\right)}^{+}\right)+2\phi_{A\left(m\right)}^{+}\right)Y_{lm}\right]\frac{1}{r}+o(r^{-1})\,.$ (30)
Note again that a $\log r/r$ term appears, arising exclusively from the advanced wave. This term, which comes both from the $1/r$ and the $1/r^{2}$ contributions to $\phi_{A}$, is time-independent and thus does not contribute to the energy flux across $\mathscr{I}^{+}$. Additionally, the advanced wave also contributes a time-independent term at order $r^{-1}$, arising again from the $1/r$ and the $1/r^{2}$ terms in $\phi_{A}$. The $\log(-u)$ term (if any) is contained in the retarded component and is at this stage unrelated to the $\log r/r$ term.
Let us now determine the asymptotic expansion near $\mathscr{I}^{-}$. By similar manipulations, one finds that the full solution reads
$\left.\phi\right|_{\mathscr{I}^{-}}$ $=-\left[\sum_{m}\phi_{R\left(m\right)}^{-}Y_{lm}\right]\frac{\log r}{r}$
$+\left[\sum_{m}\left(\frac{d}{dv}b_{m}\left(v\right)-\left(\phi_{R\left(m\right)}^{-}\log 2+\varphi_{R\left(m\right)}^{-}\right)+2\phi_{R\left(m\right)}^{-}\right)Y_{lm}\right]\frac{1}{r}+o(r^{-1})\,.$ (31)
In this case, there is a time-independent $\log r/r$ contribution arising exclusively from the retarded wave.
The expressions (30) and (31), which describe the behavior of the field near $\mathscr{I}^{\pm}$, match precisely the asymptotic expansion introduced in Ref. Fuentealba:2024lll and given in Eqs. (16) and (18), where
[…]
In the next subsection, we will show that this prescription directly yields the matching conditions for the scalar field near spatial infinity.
2.2.2 Matching conditions
Let us examine the behavior of the field $\phi$ and its canonical momentum $\pi=\partial_{t}\phi$ near spatial infinity $i^{0}$.
For the advanced solution, one must replace $v=t+r$ in the solution and then take the limit $r\rightarrow\infty$ while keeping $t$ fixed. This gives
$\left.\phi_{A}\right|_{i^{0}}$ $=\sum_{m}\frac{1}{r}\left[\phi_{A\left(m\right)}^{+}\left(\log\left(t+r\right)\right)+\varphi_{A\left(m\right)}^{+}\right]Y_{1m}$
$\qquad\qquad-\sum_{m}\frac{1}{r^{2}}\left[\phi_{A\left(m\right)}^{+}\left(\left(t+r\right)\log\left(t+r\right)-\left(t+r\right)\right)+\varphi_{A\left(m\right)}^{+}\left(t+r\right)\right]Y_{1m}+\dots\,$,
$=\left[\sum_{m}\phi_{A\left(m\right)}^{+}Y_{1m}\right]\frac{1}{r}+\dots\,.$
Note that the $\frac{\log r}{r}$ term drops.
Similarly, for the retarded solution one replaces u=tru=t-r in the solution and then takes the limit $r\rightarrow\infty$ while maintaining $t$ fixed. One gets this time
$\left.\phi_{R}\right|_{i^{0}}=\left[\sum_{m}\phi_{R\left(m\right)}^{-}Y_{1m}\right]\frac{1}{r}+\dots\,$,
again with no $\frac{\log r}{r}$ term.
Adding the two contributions, we get
$\left.\phi\right|_{i^{0}}=\left[\sum_{m}\left(\phi_{A\left(m\right)}^{+}+\phi_{R\left(m\right)}^{-}\right)Y_{1m}\right]\frac{1}{r}+\dots\,$, (34)
Let us now compute the canonical momentum $\pi=\partial_{t}\phi$ near $i^{0}$. The contribution from the advanced solution is given by
$\left.\pi_{A}\right|_{i^{0}}$ $=\partial_{t}\sum_{m}\frac{1}{r}\left[\phi_{A\left(m\right)}^{+}\left(\log\left(t+r\right)\right)+\varphi_{A\left(m\right)}^{+}\right]Y_{1m}$
$\qquad\qquad-\partial_{t}\sum_{m}\frac{1}{r^{2}}\left[\phi_{A\left(m\right)}^{+}\left(\left(t+r\right)\log\left(t+r\right)-\left(t+r\right)\right)+\varphi_{A\left(m\right)}^{+}\left(t+r\right)\right]Y_{1m}+\dots\,$,
$=-\left[\sum_{m}\phi_{A\left(m\right)}^{+}Y_{1m}\right]\frac{\log r}{r^{2}}+\left[\sum_{m}\left(\phi_{A\left(m\right)}^{+}-\varphi_{A\left(m\right)}^{+}\right)Y_{1m}\right]\frac{1}{r^{2}}+\dots\,.$
Similarly, the retarded component of the momentum near $i^{0}$ takes the form
$\left.\pi_{R}\right|_{i^{0}}$ $=\partial_{t}\sum_{m}\frac{1}{r}\left[\phi_{R\left(m\right)}^{-}\left(\log\left(r-t\right)\right)+\varphi_{R\left(m\right)}^{-}\right]Y_{1m}$
$\qquad\qquad+\partial_{t}\sum_{m}\frac{1}{r^{2}}\left[\phi_{R\left(m\right)}^{-}\left(\left(t-r\right)\log\left(r-t\right)-\left(t-r\right)\right)+\varphi_{R\left(m\right)}^{-}\left(t-r\right)\right]Y_{1m}+\dots\,$,
$=\left[\sum_{m}\phi_{R\left(m\right)}^{-}Y_{1m}\right]\frac{\log r}{r^{2}}+\left[\sum_{m}\left(-\phi_{R\left(m\right)}^{-}+\varphi_{R\left(m\right)}^{-}\right)Y_{1m}\right]\frac{1}{r^{2}}+\dots\,.$
Thus, the combined contribution of the retarded and advanced components of the momentum near $i^{0}$ becomes
$\left.\pi\right|_{i^{0}}$ $=-\left[\sum_{m}\left(\phi_{A\left(m\right)}^{+}-\phi_{R\left(m\right)}^{-}\right)Y_{1m}\right]\frac{\log r}{r^{2}}$
$-\left[\sum_{m}\left(\varphi_{A\left(m\right)}^{+}-\varphi_{R\left(m\right)}^{-}+\phi_{R\left(m\right)}^{-}-\phi_{A\left(m\right)}^{+}\right)Y_{lm}\right]\frac{1}{r^{2}}+\dots\,$ (35)
As expressed in Eqs. (7) and (8), terms of order $\log r/r$ in $\left.\phi\right|_{i^{0}}$ and $\log r/r^{2}$ in $\left.\pi\right|_{i^{0}}$ are not allowed. Therefore, from Eq. (35), we obtain the following condition between the coefficients $\phi_{A\left(m\right)}^{+}$ and $\phi_{R\left(m\right)}^{-}$
$\phi_{A\left(m\right)}^{+}=\phi_{R\left(m\right)}^{-}.$ (36)
The above expression establishes a relation between the leading logarithmic contributions of the field at $\mathscr{I}_{-}^{+}$ and $\mathscr{I}_{+}^{-}$. Using the definitions of the radial logarithmic terms given in Eqs.(32) and (33), and imposing the condition (36), one obtains:
[…]
$=0\,$
[…]
established in Fuentealba:2024lll .
If we assume that there is no leading logarithmic term at null infinity, i.e., $\phi_{A\left(m\right)}^{+}=0=\phi_{R\left(m\right)}^{-}$, then the next $1/r$ term (which becomes leading) reads
$\left.\phi\right|_{\mathscr{I}^{+}}$ $=\frac{\Phi\left(\hat{x}\right)}{r}+o(r^{-1})\,,\qquad\Phi\left(\hat{x}\right)=\sum_{m}\left(\varphi_{R\left(m\right)}^{-}-\varphi_{A\left(m\right)}^{+}\right)Y_{1m}\,$, (37)
$\left.\phi\right|_{\mathscr{I}^{-}}$ $=\frac{\Phi^{\prime}\left(\hat{x}\right)}{r}+o(r^{-1})\,,\qquad\Phi^{\prime}\left(\hat{x}\right)=\sum_{m}\left(\varphi_{A\left(m\right)}^{+}-\varphi_{R\left(m\right)}^{-}\right)Y_{1m}\,$, (38)
from which one infers the standard matching condition Campiglia:2017dpg
[…]
At the same time, the coefficient of the leading ($1/r$) term in the expansion of $\phi$ near spatial infinity vanishes while the coefficient of the leading ($1/r^{2}$) term in the expansion of $\pi$ is arbitrary, in agreement with the parity conditions recalled in the introduction (only even (respectively, odd) spherical harmonics for the leading term of $\phi$ (respectively, $\pi$)).
2.3 General case
The general case is treated exactly along the same lines as the monopole and dipole terms. As the formulas become rather involved, it is dealt with in detail in Appendix A. One finds in particular that the condition for the absence at spatial infinity of $\log r/r$ terms in $\phi$ ($l$ even) and $\log r/r^{2}$ terms in $\pi$ ($l$ odd) yields the generalized matching conditions of Fuentealba:2024lll .
One way to understand the reason that it is the asymptotic behaviour at $i^{0}$ of $\phi$ for $l$ even and that of $\pi$ for $l$ odd that give non trivial conditions goes as follows. Assume that one adopts initial values of $\phi$ and $\pi$ that violate the asymptotic decay (7)-(8) at spatial infinity by terms of the form originating from the logarithms at null infinty encountered previously. That is, consider an asymptotic behaviour that reads
$\phi|_{t=0}\sim\lambda(x^{A})\frac{\log r}{r}+\textrm{allowed terms}\,,\quad\pi|_{t=0}\sim\mu(x^{A})\frac{\log r}{r^{2}}+\textrm{allowed terms}\,$,
where the coefficients $\lambda(x^{A})$ and $\mu(x^{A})$ are a priori arbitrary. One can easily integrate the Klein-Gordon equation for such initial data in hyperbolic coordinates, all the way to null infinity, by following the method of Henneaux:2018mgn ; Fuentealba:2024lll . The differential equation that controls the time development of $\lambda(x^{A})$ and $\mu(x^{A})$ can be reduced to the Legendre differential equation. Its solution can be split into a sum of (i) a $P$-branch, which is polynomial and determined by the even part of $\lambda$ and the odd part of $\pi$; and (ii) a $Q$-branch, which exhibits a logarithmic behaviour near null infinity and which is determined by the odd part of $\lambda$ and the even part of $\mu$. The $Q$-branch is absent if and only if $\lambda^{\textrm{odd}}=0$ and $\mu^{\textrm{even}}=0$.
By integrating the equations from the initial slice to null infinity, one then finds that for arbitrary $\lambda(x^{A})$ and $\mu(x^{A})$ fulfilling no particular parity properties, the field $\phi$ takes near null infinity the form
$\phi\sim\Theta^{Q}\,\frac{(\log r)^{2}}{r}+\textrm{subleading terms}\,$,
where the coefficient $\Theta^{Q}$ vanishes if and only if the $Q$-branch is absent. In $\frac{(\log r)^{2}}{r}$, one logarithm originates from the assumed decay at spatial infinity ($\sim\frac{\log r}{r}$) while the other logarithm originates from the fact that the $Q$-branch develops a logarithmic behaviour near null infinity. The $P$-branch maintains the $\frac{\log r}{r}$ behaviour.
Now, the $\frac{(\log r)^{2}}{r}$-term is incompatible with the assumptions made at null infinity, which only allowed $\frac{\log r}{r}$-terms. From what we have seen, this means that our assumptions at null infinity automatically enforce $\lambda^{\textrm{odd}}=0$ and $\mu^{\textrm{even}}=0$. The only non-trivial components of $\lambda(x^{A})$ and $\mu(x^{A})$ present in our context are therefore $\lambda^{\textrm{even}}$ and $\mu^{\textrm{odd}}$. The requirement that the initial conditions have no $\frac{\log r}{r}$-term for $\phi$ and $\frac{\log r}{r^{2}}$-term for $\pi$ reduces accordingly to the non-trivial conditions $\lambda^{\textrm{even}}=0$ and $\mu^{\textrm{odd}}=0$ since $\lambda^{\textrm{odd}}=0$ and $\mu^{\textrm{even}}=0$ are automatically implemented [^4] Recall that with initial data that take the form (7)-(8), with no $\log$’s, the $\frac{\log r}{r}$-term at null infinity originates from the $Q$-branch of the $\frac{1}{r}$-term of the initial data Henneaux:2018mgn ; Fuentealba:2024lll ..
If we restrict our attention to a definite multipole type, $\lambda$ will thus be identically zero for $l$ odd. Requiring it to vanish will thus yield no non trivial condition. Similarly, $\mu$ will be identically zero for $l$ even so that requiring it to vanish will yield no non trivial condition.
Whatever the parity of the spherical harmonics that bring the non trivial matching conditions, we see that these originate from requirements on the asymptotic behaviour at $i^{0}$. Without these requirements, the advanced and retarded parts of the scalar field would be completely independent.
3 Electromagnetism
3.1 The electromagnetic potential
The previous analysis can be naturally extended to the case of the electromagnetic field. As in the scalar case, the last advanced wave compatible with finite incoming energy leaves a $\log r$ imprint in the field strength at $\mathscr{I}^{+}$, while the first retarded wave produces a similar imprint at $\mathscr{I}^{-}$. Consistency with the standard fall-off near $i^{0}$ then requires the matching conditions derived in Fuentealba:2025ekj .
In this section, we consider the general case that includes all possible multipole moments. As a first step, one must determine the general solution to the free Maxwell equations in retarded null coordinates.
The wave equation for the electromagnetic potential is
$\left(\boxempty A_{\mu}-\partial_{\mu}\nabla_{\nu}A^{\nu}\right)=0\,.$ (39)
Expressing the solution in retarded null coordinates, for simplicity in the gauge $A_{r}=0$, one finds the following set of equations:
$\partial_{r}^{2}A_{u}+\frac{2}{r}\partial_{r}A_{u}-\partial_{u}\partial_{r}A_{u}+\frac{1}{r^{2}}\left(D^{2}A_{u}-\partial_{u}\left(D^{B}A_{B}\right)\right)$ $=0\,$, (40)
$\partial_{r}^{2}A_{u}+\frac{2}{r}\partial_{r}A_{u}-\frac{1}{r^{2}}\partial_{r}\left(D^{B}A_{B}\right)$ $=0\,$, (41)
$\partial_{r}^{2}A_{A}-2\partial_{u}\partial_{r}A_{A}+\partial_{A}\partial_{r}A_{u}+\frac{D^{B}D_{B}A_{A}-D_{A}D_{C}A^{C}-A_{A}}{r^{2}}$ $=0\,.$ (42)
Note that, although the gauge $A_{r}=0$ has been argued in Fuentealba:2025ekj to be improper because it hides the angle-dependent logarithmic gauge transformations Fuentealba:2022xsz ; Fuentealba:2023rvf , it will be sufficient for our purposes, since we shall focus on the field strength. Furthermore, even though we work in retarded coordinates, we consider the full vector potential, describing both the retarded and the advanced waves.
The procedure for solving Eqs. (40)-(42) is described in detail in Appendix B. The full expression for the electromagnetic potential (retarded + advanced) then reads:
$A_{u}=$ $-\sum_{l,m}\sum_{k=1}^{l}\frac{1}{2^{k-1}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\frac{d^{l-k+1}}{du^{l-k+1}}g_{lm}^{\text{ret}}\left(u\right)Y_{lm}\,$,
$+\sum_{l,m}\sum_{k=0}^{l}\frac{k}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)Y_{lm}+\frac{Q}{r}\,.$
$A_{A}=$ $\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)\partial_{A}Y_{lm}$
$+2\sum_{l,m}\frac{d^{l}}{du^{l}}g_{lm}^{\text{ret}}\left(u\right)\partial_{A}Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\frac{d^{l-k}}{du^{l-k}}h_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}h_{lm}^{\text{adv}}\left(v\right)\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\,.$
The solution depends on the functions $g_{lm}^{\text{ret}}\left(u\right)$ and $g_{lm}^{\text{adv}}\left(v\right)$, associated with the “longitudinal” or “electric” sector, and $h_{lm}^{\text{ret}}\left(u\right)$ and $h_{lm}^{\text{adv}}\left(v\right)$, associated with the “transverse” or “magnetic” sector. In the electromagnetic case, there is an additional set of arbitrary functions depending on the retarded and advanced times compared to the scalar field, reflecting the fact that the electromagnetic field carries two local degrees of freedom per spatial point.
From the vector potential, we easily derive the field strength,
$F_{ur}=$ $\sum_{l,m}\sum_{k=0}^{l}\frac{k}{2^{k-1}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k+1}}{du^{l-k+1}}g_{lm}^{\text{ret}}\left(u\right)-\left(-1\right)^{k}\frac{d^{l-k+1}}{dv^{l-k+1}}g_{lm}^{\text{adv}}\left(v\right)\right)Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{k\left(k+1\right)}{2^{k}k!r^{k+2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)Y_{lm}+\frac{Q}{r^{2}}\,$,
$F_{uA}=$ $\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!\,r^{k}}\frac{(l+k)!}{(l-k)!}\Bigg(\frac{d^{\,l-k+1}}{du^{\,l-k+1}}g_{lm}^{\text{ret}}(u)+(-1)^{k}\frac{d^{\,l-k+1}}{dv^{\,l-k+1}}g_{lm}^{\text{adv}}(v)\Bigg)\partial_{A}Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{k}{2^{k}k!\,r^{k+1}}\frac{(l+k)!}{(l-k)!}\Bigg(\frac{d^{\,l-k}}{du^{\,l-k}}g_{lm}^{\text{ret}}(u)-(-1)^{k}\frac{d^{\,l-k}}{dv^{\,l-k}}g_{lm}^{\text{adv}}(v)\Bigg)\partial_{A}Y_{lm}$
$+\,\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!\,r^{k}}\frac{(l+k)!}{(l-k)!}\Big(\frac{d^{\,l-k+1}}{du^{\,l-k+1}}h_{lm}^{\text{ret}}(u)+(-1)^{k}\frac{d^{\,l-k+1}}{dv^{\,l-k+1}}h_{lm}^{\text{adv}}(v)\Big)\sqrt{\gamma}\,\epsilon_{AB}\partial^{B}Y_{lm}\,.$
$F_{rA}=$ $\sum_{l,m}\sum_{k=0}^{l}\Bigg[\frac{k}{2^{k}k!\,r^{k+1}}\frac{(l+k)!}{(l-k)!}\Bigg(\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}(u)-(-1)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}(v)\Bigg)$
$\quad\hskip 156.49014pt+\frac{2(-1)^{k}}{2^{k}k!\,r^{k}}\frac{(l+k)!}{(l-k)!}\frac{d^{l-k+1}}{dv^{l-k+1}}g_{lm}^{\text{adv}}(v)\Bigg]\partial_{A}Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\Bigg[\frac{2(-1)^{k}}{2^{k}k!\,r^{k}}\frac{(l+k)!}{(l-k)!}\frac{d^{l-k+1}}{dv^{l-k+1}}h_{lm}^{\text{adv}}(v)$
$\quad-\frac{k}{2^{k}k!\,r^{k+1}}\frac{(l+k)!}{(l-k)!}\Bigg(\frac{d^{l-k}}{du^{l-k}}h_{lm}^{\text{ret}}(u)+(-1)^{k}\frac{d^{l-k}}{dv^{l-k}}h_{lm}^{\text{adv}}(v)\Bigg)\Bigg]\sqrt{\gamma}\,\epsilon_{AB}\partial^{B}Y_{lm}\,.$
$F_{AB}=$ $\sum_{l,m}\sum_{k=0}^{l}\frac{1}{k!\,(2r)^{k}}\frac{(l+k)!}{(l-k)!}\Bigg(\frac{d^{l-k}}{du^{l-k}}h_{lm}^{\text{ret}}(u)+(-1)^{k}\frac{d^{l-k}}{dv^{l-k}}h_{lm}^{\text{adv}}(v)\Bigg)$
$\times\Big(\epsilon_{BC}\,\partial_{A}\partial^{C}Y_{lm}-\epsilon_{AC}\,\partial_{B}\partial^{C}Y_{lm}\Big)\,.$
Once the electromagnetic potential is known in retarded coordinates, one can derive its form in advanced coordinates by making the corresponding change of coordinates. However, one must also perform a change of gauge if one wants the resulting potential to fulfill the gauge conditions $A_{r}=0$, $A_{v}=O(1/r)$. Such a complication does not occur for the gauge-invariant field strength, which we shall thus simply derive in advanced coordinates from their expression in retarded coordinates through the corresponding change of variables.
3.2 Finite energy flux conditions
As in the scalar case, the functions of $u$ and $v$ appearing in the general solution are not arbitrary but constrained by finite energy flux conditions.
The energy flux is determined by the integral of the Pointing vector
$\frac{dE}{dt}=\int_{s^{2}}ds_{j}\,E^{i}F_{ij}\,.$
Let us write the above expression in spherical coordinates:
$\frac{dE}{dt}=\int_{s^{2}}d^{2}\hat{x}\,\sqrt{\gamma}F_{Ar}E^{A}\,.$
where
$F_{Ar}$ $=\partial_{A}A_{r}-\partial_{r}A_{A}\,$,
$E^{A}$ $=\gamma^{AB}\left(\partial_{A}A_{t}-\partial_{t}A_{A}\right)\,.$
Here, the components are evaluated in $(t,r,x^{A})$ coordinates. For the retarded component of the solution, one has  [^5] Using $A_{r}=0$ and $A_{u}=O\left(r^{-1}\right)$ in $(u,r,x^{A})$ coordinates implies $A_{r}=O\left(r^{-1}\right)$, $A_{t}=O\left(r^{-1}\right)$ in $(t,r,x^{A})$ coordinates
$F_{Ar}$ $=\partial_{A}A_{r}\left(t-r,\hat{x}\right)-\partial_{r}A_{A}\left(t-r,\hat{x}\right)=\partial_{u}\bar{A}_{A}\left(u,\hat{x}\right)+O\left(r^{-1}\right)\,$,
$E^{A}$ $=\gamma^{AB}\left(\partial_{A}A_{t}-\partial_{t}A_{A}\right)=-\gamma^{AB}\partial_{u}\bar{A}_{A}\left(u,\hat{x}\right)+O\left(r^{-1}\right),$
where
$A_{A}\left(u,r,\hat{x}\right)=\bar{A}_{A}\left(u,\hat{x}\right)+O\left(r^{-1}\right)$
One then gets
$\frac{dE}{du}=-\int_{s^{2}}d^{2}\hat{x}\,\sqrt{\gamma}\gamma^{AB}\left(\partial_{u}\bar{A}_{A}\right)\left(\partial_{u}\bar{A}_{B}\right)\,.$
Therefore, the total radiated energy through $\mathscr{I}^{+}$ by the retarded component of the electromagnetic field is given by
$\left.\Delta E\right|_{\mathscr{I}^{+}}=-\int_{\mathscr{I}^{+}}dud^{2}\hat{x}\,\sqrt{\gamma}\gamma^{AB}\left(\partial_{u}\bar{A}_{A}\right)\left(\partial_{u}\bar{A}_{B}\right)\,.$
Requiring this expression to be finite imposes, as in the scalar field case,
$\lim_{u\rightarrow\pm\infty}\frac{d^{l}}{du^{l}}g_{lm}^{\text{ret}}\left(u\right)$ $=g_{lm}^{\pm\text{ret}}\log\left(\pm u\right)+G_{lm}^{\pm\text{ret}}+\dots\,$, (43)
$\lim_{u\rightarrow\pm\infty}\frac{d^{l}}{du^{l}}h_{lm}^{\text{ret}}\left(u\right)$ $=h_{lm}^{\pm\text{ret}}\log\left(\pm u\right)+H_{lm}^{\pm\text{ret}}+\dots\,$, (44)
where $g_{lm}^{\text{ret}}$, $G_{lm}^{\pm\text{ret}}$, $h_{lm}^{\text{ret}}$ and $H_{lm}^{\pm\text{ret}}$ are constants. It follows by integration that
$\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)$ $\underset{u\rightarrow-\infty}{=}g_{lm}^{-\text{ret}}\left(\frac{u^{k}}{k!}\log\left(-u\right)-\frac{1}{k!}H_{k}u^{k}\right)+G_{lm}^{-\text{ret}}\frac{u^{k}}{k!}+\dots\,$,
$\frac{d^{l-k}}{du^{l-k}}h_{lm}^{\text{ret}}\left(u\right)$ $\underset{u\rightarrow-\infty}{=}h_{lm}^{-\text{ret}}\left(\frac{u^{k}}{k!}\log\left(-u\right)-\frac{1}{k!}H_{k}u^{k}\right)+H_{lm}^{-\text{ret}}\frac{u^{k}}{k!}+\dots\,.$
Here, $H_{k}=\sum_{n=1}^{k}\frac{1}{n}$ represents the harmonic numbers. The integration constants have been omitted, as they contribute only subleading terms and are therefore irrelevant for the present analysis (see formulas in appendix A).
Analogously, finiteness of the total incoming energy through $\mathscr{I}^{-}$ from the advanced field yields the conditions
$\lim_{v\rightarrow\pm\infty}\frac{d^{l}}{dv^{l}}g_{lm}^{\text{adv}}\left(v\right)$ $=g_{lm}^{\pm\text{adv}}\log\left(\pm v\right)+G_{lm}^{\pm\text{adv}}+\dots\,$, (45)
$\lim_{v\rightarrow\pm\infty}\frac{d^{l}}{dv^{l}}h_{lm}^{\text{adv}}\left(v\right)$ $=h_{lm}^{\pm\text{adv}}\log\left(\pm v\right)+H_{lm}^{\pm\text{adv}}+\dots\,$, (46)
from which one gets
$\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)$ $\underset{v\rightarrow+\infty}{=}g_{lm}^{+\text{adv}}\left(\frac{v^{k}}{k!}\log\left(v\right)-\frac{1}{k!}H_{k}v^{k}\right)+G_{lm}^{+\text{adv}}\frac{v^{k}}{k!}+\dots\,$, (47)
$\frac{d^{l-k}}{dv^{l-k}}h_{lm}^{\text{adv}}\left(v\right)$ $\underset{v\rightarrow+\infty}{=}h_{lm}^{+\text{adv}}\left(\frac{v^{k}}{k!}\log\left(v\right)-\frac{1}{k!}H_{k}v^{k}\right)+H_{lm}^{+\text{adv}}\frac{v^{k}}{k!}+\dots\,.$ (48)
3.3 Asymptotic behaviour near $i^{0}$
Further conditions arise from the requirement that the field strength involves no logarithm at spatial infinity. These relate the advanced functions to the retarded functions and will be key in the derivation of the matching conditions below.
We shall work out the components $\tilde{F}_{\mu\nu}$ of the electromagnetic tensor in $\left(t,r,x^{A}\right)$ coordinates from the components in retarded coordinates. One has
$\tilde{F}_{tr}=F_{ur}\,,\qquad\qquad\tilde{F}_{tA}=F_{uA}\,,\qquad\qquad\tilde{F}_{rA}=F_{rA}-F_{uA}\,,\qquad\qquad\tilde{F}_{AB}=F_{AB}\,.$
Since our goal is to determine the conditions that enforce the absence of $\log r$ terms at spatial infinity, we trace only the $\log u$ and $\log v$ terms in the field strength, since these are the only source of $\log r$ terms.
Let us first focus on the component $\tilde{F}_{tr}=F_{ur}$. Using the relations u=tru=t-r and $v=t+r$ as well as the expressions for the retarded and advanced functions resulting from the finite radiated/incoming energy conditions, we find (keeping only the logs)
$\tilde{F}_{tr}=\sum_{l,m}\sum_{k=0}^{l}\frac{k}{2^{k-1}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-g_{lm}^{-\text{ret}}\frac{k\left(t-r\right)^{k-1}}{k!}\log\left(r-t\right)-\right.\\ \left.\left(-1\right)^{k}g_{lm}^{+\text{adv}}k\frac{\left(t+r\right)^{k-1}}{k!}\log\left(t+r\right)\right)Y_{lm}\\ +\sum_{l,m}\sum_{k=0}^{l}\frac{k\left(k+1\right)}{2^{k}k!r^{k+2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\left(-1\right)^{k}g_{lm}^{+\text{adv}}\frac{\left(t+r\right)^{k}}{k!}\log\left(t+r\right)-g_{lm}^{-\text{ret}}\frac{\left(t-r\right)^{k}}{k!}\log\left(r-t\right)\right)Y_{lm}.$
Therefore, in the limit when $r\rightarrow\infty$, with $t$ held fixed, we find
$\left.\tilde{F}_{tr}\right|_{i^{0}}=\frac{\log r}{r^{2}}\sum_{l,m}\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k\left(k-1\right)}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\right)\left(g_{lm}^{-\text{ret}}-g_{lm}^{+\text{adv}}\right)Y_{lm}+\cdots\,.$
Expressing the sum in terms of the derivatives of the Legendre polynomials $P_{l}(x)$ evaluated at zero (see (95)) we can write
$\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k\left(k-1\right)}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}=P^{\prime\prime}_{l}(0).$ (49)
Notice that the above relation vanishes for odd values of $l$, and is non-zero for even values of $l$. Let us now compute $\tilde{F}_{tA}=F_{uA}$. Following the same rules as for $\tilde{F}_{tr}$, we find
$\tilde{F}_{tA}$ $=\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(g_{lm}^{-\text{ret}}\frac{k\left(t-r\right)^{k-1}}{k!}\log\left(r-t\right)\right$.
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left.\left(-1\right)^{k}g_{lm}^{+\text{adv}}k\frac{\left(t+r\right)^{k-1}}{k!}\log\left(t+r\right)\right)\partial_{A}Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{k}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(g_{lm}^{-\text{ret}}\frac{\left(t-r\right)^{k}}{k!}\log\left(r-t\right)\right$.
$\left.\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\left(-1\right)^{k}g_{lm}^{+\text{adv}}\frac{\left(t+r\right)^{k}}{k!}\log\left(t+r\right)\right)\partial_{A}Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(h_{lm}^{-\text{ret}}\frac{\left(t-r\right)^{k-1}}{k!}\log\left(r-t\right)\right$.
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\left.+\left(-1\right)^{k}h_{lm}^{+\text{adv}}k\frac{\left(t+r\right)^{k-1}}{k!}\log\left(t+r\right)\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}.$
Therefore, taking the limit when $r\rightarrow\infty$, keeping $t$ fixed, one finds
$\tilde{F}_{tA}$ $=\frac{\log r}{r}\left[\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(h_{lm}^{+\text{adv}}-h_{lm}^{-\text{ret}}\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\right]+\dots\,.$
Expressing the sum in terms of the derivatives of the Legendre polynomials evaluated at zero (see equation (94) for a derivation),
$\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}=-2P^{\prime}_{l}(0)\,$, (50)
which now vanishes for even values of $l$, we obtain
$\left.\tilde{F}_{tA}\right|_{i^{0}}=\frac{\log r}{r}\left[-2\sum_{l,m}P^{\prime}_{l}(0)\left(h_{lm}^{+\text{adv}}-h_{lm}^{-\text{ret}}\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\right]+\dots\,.$
Similar considerations yield $\tilde{F}_{rA}=F_{rA}-F_{uA}$,
$\tilde{F}_{rA}=\sum_{l,m}\Bigg[\sum_{k=0}^{l}\frac{1}{2^{k}k!\,r^{k}}\frac{(l+k)!}{(l-k)!}\Bigg(-g_{lm}^{-\text{ret}}\frac{k\,(t-r)^{k-1}}{k!}\log(r-t)$
$\hskip-113.81102pt+(-1)^{k}g_{lm}^{+\text{adv}}\frac{k\,(t+r)^{k-1}}{k!}\log(t+r)\Bigg)\Bigg]\partial_{A}Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!\,r^{k}}\frac{(l+k)!}{(l-k)!}\Bigg(-h_{lm}^{-\text{ret}}\frac{k\,(t-r)^{k-1}}{k!}\log(r-t)$
$\hskip-113.81102pt+(-1)^{k}h_{lm}^{+\text{adv}}\frac{k\,(t+r)^{k-1}}{k!}\log(t+r)\Bigg)\sqrt{\gamma}\,\epsilon_{AB}\,\partial^{B}Y_{lm}$
$-\sum_{l,m}\Bigg[\sum_{k=0}^{l}\frac{k}{2^{k}k!\,r^{k+1}}\frac{(l+k)!}{(l-k)!}\Bigg(h_{lm}^{-\text{ret}}\frac{(t-r)^{k}}{k!}\log(r-t)$
$\hskip-113.81102pt+(-1)^{k}h_{lm}^{+\text{adv}}\frac{(t+r)^{k}}{k!}\log(t+r)\Bigg)\Bigg]\sqrt{\gamma}\,\epsilon_{AB}\,\partial^{B}Y_{lm}\,.$
After using (50), and taking the limit $r\rightarrow\infty$ with $t$ kept constant, reduces to
$\left.\tilde{F}_{rA}\right|_{i^{0}}$ $=\frac{\log r}{r}\sum_{l,m}\left[-2P_{l}^{\prime}(0)\left(g_{lm}^{-\text{ret}}+g_{lm}^{+\text{adv}}\right)\right]\partial_{A}Y_{lm}+\dots\,.$
Finally, let us turn to $\tilde{F}_{AB}=F_{AB}$. We find this time by performing similar manipulations
$F_{AB}=\sum_{l,m}\sum_{k=0}^{l}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(h_{lm}^{-\text{ret}}\frac{\left(t-r\right)^{k}}{k!^{2}\left(2r\right)^{k}}\log\left(r-t\right)+h_{lm}^{+\text{adv}}\frac{\left(-1\right)^{k}\left(t+r\right)^{k}}{k!^{2}\left(2r\right)^{k}}\log\left(t+r\right)\right)\\ \times\left(\epsilon_{BC}\partial_{A}\partial^{C}Y_{lm}-\epsilon_{AC}\partial_{B}\partial^{C}Y_{lm}\right).$
Therefore, in the limit when $r\rightarrow\infty$ keeping $t$ constant, we obtain
$F_{AB}=\log r\sum_{l,m}\left[P_{l}(0)\left(h_{lm}^{-\text{ret}}+h_{lm}^{+\text{adv}}\right)\right]\left(\epsilon_{BC}\partial_{A}\partial^{C}Y_{lm}-\epsilon_{AC}\partial_{B}\partial^{C}Y_{lm}\right),$
where we used the relation (93)
$\sum_{k=0}^{l}\frac{(-1)^{k}}{(k!)^{2}}\frac{(l+k)!}{(l-k)!}=P_{l}(0).$ (51)
In order to eliminate the leading $\log r$ terms, we must impose the following requirements
$P_{l}^{\prime}(0)\left(h_{lm}^{+\text{adv}}-h_{lm}^{-\text{ret}}\right)=0,\quad P_{l}(0)\left(h_{lm}^{-\text{ret}}+h_{lm}^{+\text{adv}}\right)=0,$ (52)
and
$P^{\prime}_{l}(0)\left(g_{lm}^{-\text{ret}}+g_{lm}^{+\text{adv}}\right)=0,\quad P^{\prime\prime}_{l}(0)\left(g_{lm}^{-\text{ret}}-g_{lm}^{+\text{adv}}\right)=0.$ (53)
These have the same form as in the scalar case discussed in (72) and (73). They lead to the following conditions,
$g_{lm}^{-\text{ret}}=\left(-1\right)^{l}g_{lm}^{+\text{adv}}\,,\qquad\qquad h_{lm}^{-\text{ret}}=-\left(-1\right)^{l}h_{lm}^{+\text{adv}}\,.$ (54)
These relations between the advanced and retarded fields are essential for establishing the matching conditions connecting past and future null infinity for the logarithmic terms.
Note that, as in the scalar field case, the contributions to the leading logarithmic term at $i^{0}$ arise from all powers of $r$ appearing in the electromagnetic tensor expressed in retarded or advanced coordinates. Accordingly, it becomes necessary to extend the near-$\mathcal{I^{\pm}}$ expansion into the bulk in order to derive the conditions (54) and determine the leading orders in the vicinity of $i^{0}$.
3.4 Matching conditions
The matching conditions of He:2014cra ; Kapec:2015ena ; Strominger:2017zoo relate the leading order of the radial electric and magnetic fields at the past of $\mathscr{I}^{+}$ to the leading order of the corresponding fields at the future of $\mathscr{I}^{-}$. They were derived assuming no logarithmic terms at null infinity. To derive them in the general case, let us thus examine the behaviour of the fields near null infinity, starting with the electromagnetic potential.
3.4.1 Expansion of the electromagnetic potential near $\mathscr{I}^{+}$
The advanced solution is by definition the $v$-dependent part of the electromagnetic potential, i.e.,
$A_{u}^{\text{adv}}=$ $\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)Y_{lm}\,$,
$A_{A}^{\text{adv}}=$ $\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)\partial_{A}Y_{lm}$
$+\sum_{l,l}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\frac{d^{l-k}}{dv^{l-k}}h_{lm}^{\text{adv}}\left(v\right)\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\,.$
When expanded near future null infinity, it brings $\log r$ terms, as in the scalar case, which we now determine.
Using the expressions (45)-(48) for $g_{lm}^{\text{adv}}$ and $h_{lm}^{\text{adv}}$, the advanced solution becomes
$A_{u}^{\text{adv}}$ $=\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{2^{k}k!^{2}r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[g_{lm}^{+\text{adv}}\left(v^{k}\log\left(v\right)-H_{k}v^{k}\right)+G_{lm}^{+\text{adv}}v^{k}+\dots\right]Y_{lm}\,$,
$A_{A}^{\text{adv}}$ $=\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}k!^{2}r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[g_{lm}^{+\text{adv}}\left(v^{k}\log\left(v\right)-H_{k}v^{k}\right)+G_{lm}^{+\text{adv}}v^{k}+\dots\right]\partial_{A}Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}k!^{2}r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[h_{lm}^{+\text{adv}}\left(v^{k}\log\left(v\right)-H_{k}v^{k}\right)+H_{lm}^{+\text{adv}}v^{k}+\dots\right]\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}.$
Substituting $v=u+2r$, and expanding for large values of $r$, i.e., near future null infinity, we find
$A_{u}^{\text{adv}}$ $=\frac{\log r}{r}\left[\sum_{l,m}g_{lm}^{+\text{adv}}\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\right)Y_{lm}\right]+$
$\,\,\,\,\,\,\,\,\,\,\,\frac{1}{r}\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[g_{lm}^{+\text{adv}}\log 2-g_{lm}^{+\text{adv}}H_{k}+G_{lm}^{+\text{adv}}\right]Y_{lm}+\dots\,$,
$A_{A}^{\text{adv}}$ $=\log r\,\sum_{l,m}\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\right)\left[g_{lm}^{+\text{adv}}\partial_{A}Y_{lm}+h_{lm}^{+\text{adv}}\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\right]$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[g_{lm}^{+\text{adv}}\log 2-g_{lm}^{+\text{adv}}H_{k}+G_{lm}^{+\text{adv}}\right]\partial_{A}Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[h_{lm}^{+\text{adv}}\log 2-h_{lm}^{+\text{adv}}H_{k}+H_{lm}^{+\text{adv}}\right]\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}$
$+\frac{\log r}{r}\left[\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{2\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}u\left(g_{lm}^{+\text{adv}}\partial_{A}Y_{lm}+h_{lm}^{+\text{adv}}\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\right)\right]+\dots\,.$
The above expressions can be simplified using the identities (96) and (97)
$\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}=\left(-1\right)^{l}\,,\qquad\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}=\left(-1\right)^{l}l\left(l+1\right)\,$, (55)
which imply
$\left.A_{u}^{\text{adv}}\right|_{\mathscr{I}^{+}}$ $=\frac{\log r}{r}\left[\sum_{l,m}\left(-1\right)^{l}l\left(l+1\right)g_{lm}^{+\text{adv}}Y_{lm}\right]+$
$\frac{1}{r}\sum_{l,m}\left[\left(-1\right)^{l}l\left(l+1\right)\left(g_{lm}^{+\text{adv}}\log 2+G_{lm}^{+\text{adv}}\right)-\sum_{k=0}^{l}g_{lm}^{+\text{adv}}\frac{\left(-1\right)^{k}k}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right]Y_{lm}\,$,
$\left.A_{A}^{\text{adv}}\right|_{\mathscr{I}^{+}}$ $=\log r\sum_{l,m}\left[\left(-1\right)^{l}g_{lm}^{+\text{adv}}\partial_{A}Y_{lm}+\left(-1\right)^{l}h_{lm}^{+\text{adv}}\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\right]$
$+\sum_{l,m}\left[\left(-1\right)^{l}\left(g_{lm}^{+\text{adv}}\log 2+G_{lm}^{+\text{adv}}\right)-g_{lm}^{+\text{adv}}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right]\partial_{A}Y_{lm}+$
$+\sum_{l,m}\left[\left(-1\right)^{l}\left(h_{lm}^{+\text{adv}}\log 2+H_{lm}^{+\text{adv}}\right)-h_{lm}^{+\text{adv}}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right]\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}$
$+\frac{\log r}{r}\left[\sum_{l,m}\frac{u}{2}\left(-1\right)^{l}l\left(l+1\right)\left(g_{lm}^{+\text{adv}}\partial_{A}Y_{lm}+h_{lm}^{+\text{adv}}\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\right)\right]+\dots\,.$
Therefore, the full solution (retarded + advanced) near $\mathscr{I}^{+}$ is given by
$A_{u}$ $=\frac{\log r}{r}\left[\sum_{l,m}\left(-1\right)^{l}l\left(l+1\right)g_{lm}^{+\text{adv}}Y_{lm}\right]+$
$\frac{1}{r}\sum_{l,m}\left[\left(-1\right)^{l}l\left(l+1\right)\left(g_{lm}^{+\text{adv}}\log 2+G_{lm}^{+\text{adv}}\right)-g_{lm}^{+\text{adv}}\sum_{k=0}^{l}\left(\frac{\left(-1\right)^{k}k}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right)\right]$ (56)
$\left.-\sum_{l,m}\sum_{k=1}^{l}l\left(l+1\right)\left(\frac{d^{l}}{du^{l}}g_{lm}^{\text{ret}}\left(u\right)\right)+Q\right]Y_{lm}+\dots\,$,
and
$A_{A}$ $=\log r\Bigg[\sum_{l,m}(-1)^{l}g_{lm}^{+\text{adv}}\,\partial_{A}Y_{lm}+\sum_{l,m}(-1)^{l}h_{lm}^{+\text{adv}}\,\sqrt{\gamma}\,\epsilon_{AB}\,\partial^{B}Y_{lm}\Bigg]$
$\quad+\sum_{l,m}\Bigg[\frac{d^{l}}{du^{l}}\,g_{lm}^{\text{ret}}(u)+(-1)^{l}\!\left(g_{lm}^{+\text{adv}}\log 2+G_{lm}^{+\text{adv}}\right)$
$\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,-g_{lm}^{+\text{adv}}\!\sum_{k=0}^{l}\frac{(-1)^{k}}{(k!)^{2}}\frac{(l+k)!}{(l-k)!}\,H_{k}\Bigg]\partial_{A}Y_{lm}$ (57)
$\quad+\sum_{l,m}\Bigg[\frac{d^{l}}{du^{l}}\,h_{lm}^{\text{ret}}(u)+(-1)^{l}\!\left(h_{lm}^{+\text{adv}}\log 2+H_{lm}^{+\text{adv}}\right)$
$\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,-h_{lm}^{+\text{adv}}\!\sum_{k=0}^{l}\frac{(-1)^{k}}{(k!)^{2}}\frac{(l+k)!}{(l-k)!}\,H_{k}\Bigg]\sqrt{\gamma}\,\epsilon_{AB}\,\partial^{B}Y_{lm}$
$\quad+\frac{\log r}{r}\Bigg[\sum_{l,m}\frac{u}{2}(-1)^{l}l(l+1)\big(g_{lm}^{+\text{adv}}\,\partial_{A}Y_{lm}+h_{lm}^{+\text{adv}}\,\sqrt{\gamma}\,\epsilon_{AB}\,\partial^{B}Y_{lm}\big)\Bigg]+\dots\,.$
3.4.2 Electromagnetic tensor near $\mathscr{I}^{+}$
Let us compute the electromagnetic tensor near $\mathscr{I}^{+}$, focusing on the radial electric field $F_{ur}$ and the radial magnetic field $F_{AB}$ occurring in the matching conditions of He:2014cra ; Kapec:2015ena ; Strominger:2017zoo . These components are scalars under the changes of coordinates $(u,r)\leftrightarrow(t,r)\leftrightarrow(v,r)$. We can also derive the matching conditions for the other components since we have all the expressions, but because there is nothing we can compare them with in the literature, we shall not do so here.
From the asymptotic developments of the vector potential in Eqs. (56) and (57), one gets
$F_{ur}$ $=\frac{\log r}{r^{2}}\left[\sum_{l,m}\left(-1\right)^{l}l\left(l+1\right)g_{lm}^{+\text{adv}}Y_{lm}\right]$
$+\frac{1}{r^{2}}\sum_{l,m}\left[\left(-1\right)^{l}l\left(l+1\right)(g_{lm}^{+\text{adv}}\left(\log 2-1\right)+G_{lm}^{+\text{adv}})-\sum_{k=0}^{l}g_{lm}^{+\text{adv}}\frac{\left(-1\right)^{k}H_{k}k\left(l+k\right)!}{\left(k!\right)^{2}\left(l-k\right)!}\right$.
$\left.-\sum_{k=1}^{l}l\left(l+1\right)\left(\frac{d^{l}}{du^{l}}g_{lm}^{\text{ret}}\left(u\right)\right)\right]Y_{lm}+\frac{Q}{r^{2}}+\dots\,$, (58)
and
$F_{AB}$ $=\log r\sum_{l,m}\left(-1\right)^{l}h_{lm}^{+\text{adv}}\sqrt{\gamma}\left(\epsilon_{BC}\partial_{A}\partial^{C}Y_{lm}-\epsilon_{AC}\partial_{B}\partial^{C}Y_{lm}\right)$
$+\sum_{l,m}\left[\frac{d^{l}}{du^{l}}h_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{l}\left(h_{lm}^{+\text{adv}}\log 2+H_{lm}^{+\text{adv}}\right)-h_{lm}^{+\text{adv}}\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right)\right]$
$\times\sqrt{\gamma}\left(\epsilon_{BC}\partial_{A}\partial^{C}Y_{lm}-\epsilon_{AC}\partial_{B}\partial^{C}Y_{lm}\right)+\dots\,.$
In both components of the field strength, there is a leading radial logarithmic contribution arising from the advanced wave. In particular, the leading term of $F_{ur}$ is entirely determined by the “electric” mode $g_{lm}^{+\text{adv}}$ of the advanced wave, while the leading term of $F_{AB}$ depends only on the “magnetic” advanced mode $h_{lm}^{+\text{adv}}$ of the electromagnetic wave.
3.4.3 Electromagnetic tensor near $\mathscr{I}^{-}$
Similar developments, in which one inserts u=v2ru=v-2r into the retarded solution, yield the asymptotic expansion of the field strength near $\mathscr{I}^{-}$,
$F_{vr}$ $=\frac{\log r}{r^{2}}\sum_{l,m}\left(-1\right)^{l+1}l\left(l+1\right)g_{lm}^{-\text{ret}}Y_{lm}$
$-\frac{1}{r^{2}}\sum_{l,m}\left[-l\left(l+1\right)\frac{d^{l}}{dv^{l}}g_{lm}^{\text{adv}}\left(v\right)Y_{lm}+\left(-1\right)^{l}l\left(l+1\right)\left(g_{lm}^{-\text{ret}}\left(\log 2-1\right)+G_{lm}^{-\text{ret}}\right)\right$. (59)
$\left.-\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right)g_{lm}^{-\text{ret}}Y_{lm}\right]+\dots\,$,
and
$F_{AB}$ $=\log r\sum_{l,m}\left(-1\right)^{l}h_{lm}^{-\text{ret}}\sqrt{\gamma}\left(\epsilon_{BC}\partial_{A}\partial^{C}Y_{lm}-\epsilon_{AC}\partial_{B}\partial^{C}Y_{lm}\right)$
$+\sum_{l,m}\left[\frac{d^{l}}{dv^{l}}h_{lm}^{\text{adv}}\left(v\right)+\left(-1\right)^{l}\left(h_{lm}^{-\text{ret}}\log 2+H_{lm}^{-\text{ret}}\right)-\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right)h_{lm}^{-\text{ret}}\right]$
$\times\sqrt{\gamma}\left(\epsilon_{BC}\partial_{A}\partial^{C}Y_{lm}-\epsilon_{AC}\partial_{B}\partial^{C}Y_{lm}\right)+\dots\,.$
3.4.4 Matching conditions
The matching conditions on the field strength follow immediately from the conditions relating the advanced and retarded fields obtained at $i^{0}$ in Eq. (54). Indeed, from Eqs. (58) and (59), the leading logarithmic terms of $F_{ur}$ and $F_{vr}$ are given
[…]
Therefore, using (54), we find the following matching condition:
[…]
Similarly,
[…]
and
[…]
Thus,
[…]
In this case, we must take into account that the magnetic sector has an extra minus sign in the parity coming from the Levi-Civita symbol, i.e.,
[…]
As shown in Fuentealba:2025ekj , the matching conditions for the coefficients of the leading (logarithmic) terms have opposite sign compared with those of He:2014cra ; Kapec:2015ena ; Strominger:2017zoo . If one assumes that the leading logarithmic terms are absent, the matching conditions for the coefficients of the next terms, which become leading, obey the matching conditions with the sign of He:2014cra ; Kapec:2015ena ; Strominger:2017zoo characteristic of the $P$-branch. The derivation proceeds as in the scalar case.
4 Conclusions
In this work, we have studied the role of advanced and retarded radiation in determining the asymptotic structure of massless scalar and electromagnetic fields that exhibit radial logarithmic terms near $\mathscr{I}^{\pm}$. Assuming that the total incoming energy flux across $\mathscr{I}^{-}$ is finite, we have shown that the final advanced wave, characterized by an asymptotic $\log v$ dependence at late times, leaves a distinctive imprint at $\mathscr{I}^{+}$ in the form of $\log r$ terms. Conversely, the first retarded wave completely determines the radial logarithmic behavior in the vicinity of $\mathscr{I}^{-}$. This analysis offers both a novel perspective and a coherent physical interpretation of the origin of the radial logarithmic terms. Furthermore, the results obtained through this approach coincide with those in Refs. Fuentealba:2024lll ; Fuentealba:2025ekj , where a spatial-infinity-based route was used.
Remarkably, this advanced/retarded wave approach enables the determination of the matching conditions for the radial logarithmic terms between the future of past null infinity and the past of future null infinity. This is achieved by evaluating the fields in the vicinity of $i^{0}$ and imposing that their asymptotic behaviour be of the Coulombic type, free of $\log r$ contributions at spatial infinity. The matching conditions derived using this approach are in exact agreement with those in Refs. Fuentealba:2024lll ; Fuentealba:2025ekj .
As we just stressed, the assumption that there is no logarithmic term at spatial infinity plays a key role in our derivation of the matching conditions. This assumption is motivated by the asymptotic behaviour of the elementary solution of the Poissson equation (and of superpositions of such solutions). However, this assumption can be relaxed while preserving finiteness of Poincaré charges and of the energy fluxes through past and future infinity. For instance $\log r/r$ terms with coefficients having definite parity at spatial infinity could be contemplated. It is beyond the scope of this paper to investigate the most liberal asymptotic behaviour at spatial infinity that is consistent with the above physical requirements, since the assumed $1/r$ behaviour is sufficient for our purposes.
It is natural to ask to what extent these results can be extended to the gravitational case. For linearized gravity, one may naturally expect a behaviour closely analogous to that of the scalar and electromagnetic fields discussed in this work. However, in the fully nonlinear theory, the situation is considerably more subtle. This is because, as observed for the scalar and electromagnetic fields discussed above, all subleading terms in the radial expansion near $\mathscr{I}^{-}$ contribute to the leading $\log r$ term at $\mathscr{I}^{+}$. In General Relativity, it is extremely challenging to control all subleading terms in the radial expansion due to the inherent nonlinearities of the theory. However, since Bondi’s energy-loss formula Bondi:1960jsa has a similar form to that of the scalar and electromagnetic cases, the requirement of finite total energy implies that the shear tensor at late advanced times must take the form $C_{AB}\underset{v\rightarrow\infty}{=}C_{AB}^{+}\log v+\dots$. Consequently, when it is reexpressed in retarded coordinates, it will generically give rise to $\log r$ contributions in the asymptotic expansion near $\mathscr{I}^{+}$. How to connect these logarithmic terms to the polyhomogeneous solutions involving dominant logarithmic terms presented for example, in Refs. Winic:1985 ; Chrusciel:1993hx ; Andersson:1994ng remains an interesting question for future investigation.
Acknowledgements
The research of AP is partially supported by Fondecyt grant 1220910. AP acknowledges support from the Erwin Schrödinger International Institute for Mathematics and Physics (ESI) where some of the research was undertaken during the ESI Research in Teams Programme. The work of HG and MB is funded by FONDECYT grant 1230853. The work of MH is partially supported by FNRS-Belgium (convention IISN 4.4503.15), as well as by research funds from the Solvay Family.
Appendices
Appendix A General analysis of the massless scalar field
A.1 Retarded and advanced waves
In this section, we extend the scalar field analysis of spherical and dipole waves to the general solution of the wave equation. As for the dipole case, subleading terms in the asymptotic radial expansion must be taken into account, as they play a fundamental role.
The general solution of the wave equation in a source-free region is expressed again as the sum of retarded and advanced terms.
$\phi=\phi_{R}+\phi_{A},$
with
$\phi_{R}$ $=\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\frac{d^{l-k}}{du^{l-k}}a_{l,m}\left(u\right)Y_{lm}\left(\theta,\phi\right),$ (60)
$\phi_{A}$ $=\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\frac{d^{l-k}}{dv^{l-k}}b_{l,m}\left(v\right)Y_{lm}\left(\theta,\phi\right).$ (61)
Here, $a_{l,m}\left(u\right)$ and $b_{l,m}\left(v\right)$ are arbitrary functions of the retarded and advanced times, respectively. It is noteworthy that the solution above is expressed purely as a power-law expansion in the radial coordinate, with no logarithmic terms.
In particular, the leading-order term at large $r$ takes the form:
$\phi_{R}=\frac{\bar{\phi}_{R}\left(u,\hat{x}\right)}{r}+O\left(r^{-2}\right)\,,\quad\phi_{A}=\frac{\bar{\phi}_{A}\left(v,\hat{x}\right)}{r}+O\left(r^{-2}\right)\,$,
where
[…]
The total energy flux across $\mathscr{I}^{\pm}$ takes exactly the same form as in Eqs. (10) and (12). Consequently, for the $l$-th spherical harmonic, the requirement of finite energy flux across $\mathscr{I}^{\pm}$ implies that
$\lim_{u\rightarrow\pm\infty}\frac{d^{l}}{du^{l}}a_{l,m}\left(u\right)$ $=\phi_{R\left(l,m\right)}^{\pm}\log\left(\pm u\right)+\varphi_{R\left(l,m\right)}^{\pm}+\dots\,$,
$\lim_{v\rightarrow\pm\infty}\frac{d^{l}}{dv^{l}}b_{l,m}\left(v\right)$ $=\phi_{A\left(l,m\right)}^{\pm}\log\left(\pm v\right)+\varphi_{A\left(l,m\right)}^{\pm}+\dots\,$,
assuming again integer power law behaviour for $\partial_{u}\bar{\phi}_{R}$ ($\partial_{v}\bar{\phi}_{A}$).
Let us examine the contribution of the advanced wave to the asymptotic behavior of the field near $\mathscr{I}^{+}$. This requires integrating the following equation
$\frac{d^{l}}{dv^{l}}b_{l,m}\left(v\right)\underset{v\rightarrow+\infty}{=}\phi_{A\left(l,m\right)}^{+}\log\left(v\right)+\varphi_{A\left(l,m\right)}^{+}+\dots\,.$
To carry out the integration, we will make use of the following integral identity
$\alpha\underbrace{\int\cdots\int}_{k}\log x\,dx+\beta\underbrace{\int\cdots\int}_{k}\,dx=\alpha\left(\frac{x^{k}}{k!}\log\left(x\right)-\frac{1}{k!}H_{k}x^{k}\right)+\beta\frac{x^{k}}{k!}+\sum_{n=0}^{k-1}c_{k-n}\frac{x^{n}}{n!}\,.$ (62)
Here, $c_{k-n}$ denotes the integration constants, and $H_{k}=\sum_{n=1}^{k}\frac{1}{n}$ represents the harmonic numbers. This identity can be readily proven by induction, starting from its differential form,
$\frac{d^{k}}{dx^{k}}\left[\alpha\left(\frac{x^{k}}{k!}\log\left(x\right)-\frac{1}{k!}H_{k}x^{k}\right)+\beta\frac{x^{k}}{k!}+\sum_{n=0}^{k-1}c_{k-n}\frac{x^{n}}{n!}\right]=\alpha\log\left(x\right)+\beta\,$,
and using the recursive property of the harmonic numbers, $H_{k+1}=H_{k}+\frac{1}{k+1}$.
Thus,
$\frac{d^{l-k}}{dv^{l-k}}b_{l,m}\left(v\right)\underset{v\rightarrow\infty}{=}$ $\phi_{A\left(l,m\right)}^{+}\left(\frac{v^{k}}{k!}\log\left(v\right)-\frac{1}{k!}H_{k}v^{k}\right)+\varphi_{A\left(l,m\right)}\frac{v^{k}}{k!}+\sum_{n=0}^{k-1}c_{A\left(k-n\right)}^{\left(l,m\right)}\frac{v^{n}}{n!}\,.$ (63)
The term with $k=0$ in the above expression can be properly defined by setting $H_{0}=0$. Substituting Eq. (63) into Eq. (61) one finds that the advanced solution at large $v$ can then be written as
$\phi_{A}\underset{v\rightarrow\infty}{=}\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[\phi_{A\left(l,m\right)}^{+}\left(\frac{v^{k}}{k!}\log\left(v\right)-\frac{1}{k!}H_{k}v^{k}\right)+\varphi_{A\left(l,m\right)}\frac{v^{k}}{k!}+\dots\right]Y_{lm}.$ (64)
We omit the terms containing the integration constants $c_{A(k-n)}^{(l,m)}$, as they contribute only subleading corrections that do not affect the present analysis.
The key step is to express the advanced coordinate in terms of the retarded one using $v=u+2r$. This allows one to write
$\phi_{A}=\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[\phi_{A\left(l,m\right)}^{+}\left(\frac{\left(u+2r\right)^{k}}{k!}\log\left(u+2r\right)-\frac{1}{k!}H_{k}\left(u+2r\right)^{k}\right)\right.\\ \left.+\varphi_{A\left(l,m\right)}\frac{\left(u+2r\right)^{k}}{k!}+\dots\right]Y_{lm}.$
In the limit $r\rightarrow\infty$ with $u$ held constant, the leading terms are given by
$\phi_{A}=\left[\sum_{l,m}\phi_{A\left(l,m\right)}^{+}\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\right)Y_{lm}\right]\frac{\log r}{r}\\ +\left[\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[\phi_{A\left(l,m\right)}^{+}\left(\log 2-H_{k}\right)+\varphi_{A\left(l,m\right)}\right]Y_{lm}\right]\frac{1}{r}+\dots\,.$
The above expression can be simplified using the identity (96)
$\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}=\left(-1\right)^{l}\,.$
Consequently, if one considers the sum of the advanced and retarded solutions, the full solution near $\mathscr{I}^{+}$ becomes:
$\left.\phi\right|_{\mathscr{I}^{+}}$ $=\left[\sum_{l,m}\left(-1\right)^{l}\phi_{A\left(l,m\right)}^{+}Y_{lm}\right]\frac{\log r}{r}$
[…]
$-\left.\left.\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right)\phi_{A\left(l,m\right)}^{+}\right)Y_{lm}\right]\frac{1}{r}+\dots\,.$
Note that a $\log r/r$ term appears, arising exclusively from the advanced wave. This term is time-independent and typically includes all spherical harmonics. Due to its time-independence, it does not contribute to the energy flux across $\mathscr{I}^{+}$. Additionally, the advanced wave also contributes a time-independent term at order $r^{-1}$.
Let us now determine the contribution of the retarded wave to the asymptotic expansion near $\mathscr{I}^{-}$. The requirement of finite total radiated energy implies that
$\frac{d^{l}}{du^{l}}a_{l,m}\left(u\right)\underset{u\rightarrow-\infty}{=}\phi_{R\left(l,m\right)}^{-}\log\left(-u\right)+\varphi_{R\left(l,m\right)}^{-}+\dots\,.$
Integrating the previous equation using the identity along the same lines of (62)
$\alpha\underbrace{\int\cdots\int}_{k}\log\left(-x\right)\,dx+\beta\underbrace{\int\cdots\int}_{k}\,dx=\alpha\left(\frac{x^{k}}{k!}\log\left(-x\right)-\frac{1}{k!}H_{k}x^{k}\right)+\beta\frac{x^{k}}{k!}+\sum_{n=0}^{k-1}c_{k-n}\frac{x^{n}}{n!}\,$,
one obtains
$\frac{d^{l-k}}{du^{l-k}}a_{l,m}\left(u\right)\underset{v\rightarrow-\infty}{=}$ $\phi_{R\left(l,m\right)}^{-}\left(\frac{u^{k}}{k!}\log\left(-u\right)-\frac{1}{k!}H_{k}u^{k}\right)+\varphi_{R\left(l,m\right)}^{-}\frac{u^{k}}{k!}+\sum_{n=0}^{k-1}c_{R\left(k-n\right)}^{\left(l,m\right)}\frac{u^{n}}{n!}\,.$
Substituting this result into Eq. (60), one finds the following expression for the retarded solution in the limit $u\rightarrow$-$\infty$:
$\phi_{R}=\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[\phi_{R\left(l,m\right)}^{-}\left(\frac{u^{k}}{k!}\log\left(-u\right)-\frac{1}{k!}H_{k}u^{k}\right)+\varphi_{R\left(l,m\right)}^{-}\frac{u^{k}}{k!}+\dots\right]Y_{lm}.$ (66)
By expressing the retarded time $u$ in terms of the advanced time $v$ via the relation u=v2ru=v-2r, one finds
$\phi_{R}=\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[\phi_{R\left(l,m\right)}^{-}\left(\frac{\left(v-2r\right)^{k}}{k!}\log\left(2r-v\right)-\frac{1}{k!}H_{k}\left(v-2r\right)^{k}\right)\right$.
$\left.+\varphi_{R\left(l,m\right)}^{-}\frac{\left(v-2r\right)^{k}}{k!}+\dots\right]Y_{lm}.$
Therefore, in the limit $r\rightarrow\infty$ with $v$ held constant, the leading terms are:
$\phi_{R}=$ $\left[\sum_{l,m}\phi_{R\left(l,m\right)}^{-}\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\right)Y_{lm}\right]\frac{\log r}{r}$
$+\left[\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left[\phi_{R\left(l,m\right)}^{-}\left(\log 2-H_{k}\right)+\varphi_{R\left(l,m\right)}^{-}\right]Y_{lm}\right]\frac{1}{r}+\cdots\,.$
Consequently, the full solution near $\mathscr{I}^{-}$ becomes
$\left.\phi\right|_{\mathscr{I}^{-}}$ $=\left[\sum_{l,m}\left(-1\right)^{l}\phi_{R\left(l,m\right)}^{-}Y_{lm}\right]\frac{\log r}{r}$
[…]
$\left.\left.-\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\right)\phi_{R\left(l,m\right)}^{-}Y_{lm}\right)\right]\frac{1}{r}+\cdots\,.$
In this case, there is a time-independent $\log r/r$ contribution arising exclusively from the retarded wave.
The expressions (65) and (67), which describe the behavior of the field near $\mathscr{I}^{\pm}$, match precisely the asymptotic expansion introduced in Ref. Fuentealba:2024lll and given in Eqs. (16) and (18), where
[…]
In the next subsection, we will show that this prescription directly yields the matching conditions for the scalar field near spatial infinity.
A.2 Matching conditions
Let us examine the behavior of the field $\phi$ and its canonical momentum $\pi=\partial_{t}\phi$ near spatial infinity $i^{0}$.
For the advanced solution, one must replace $v=t+r$ in Eq. (64) and then take the limit $r\rightarrow\infty$ while keeping $t$ fixed.
$\left.\phi_{A}\right|_{i^{0}}=$ $\sum_{l,m}\sum_{k=0}^{l}\frac{(-1)^{k}}{2^{k}k!\,r^{k+1}}\frac{(l+k)!}{(l-k)!}\Bigg[\phi_{A(l,m)}^{+}\!\left(\frac{(t+r)^{k}}{k!}\log(t+r)-\frac{H_{k}}{k!}(t+r)^{k}\right)$
$+\varphi_{A(l,m)}\,\frac{(t+r)^{k}}{k!}\Bigg]Y_{lm}+\dots$
$=$ $\Bigg[\sum_{l,m}\phi_{A(l,m)}^{+}\!\left(\sum_{k=0}^{l}\frac{(-1)^{k}}{2^{k}(k!)^{2}}\frac{(l+k)!}{(l-k)!}\right)Y_{lm}\Bigg]\frac{\log r}{r}$
$+\Bigg[\sum_{l,m}\sum_{k=0}^{l}\frac{(-1)^{k}}{2^{k}(k!)^{2}}\frac{(l+k)!}{(l-k)!}\big(\varphi_{A(l,m)}-\phi_{A(l,m)}^{+}H_{k}\big)Y_{lm}\Bigg]\frac{1}{r}+\dots\,.$
Similarly, for the retarded solution one replaces u=tru=t-r in Eq. (66) and then take the limit $r\rightarrow\infty$ while maintaining $t$ fixed.
$\left.\phi_{R}\right|_{i^{0}}$ $=\sum_{l,m,k\leq l}\frac{1}{2^{k}k!\,r^{k+1}}\frac{(l+k)!}{(l-k)!}\frac{(t-r)^{k}}{k!}\Bigg(\phi_{R(l,m)}^{-}\!\left(\log(r-t)-H_{k}\right)+\varphi_{R(l,m)}^{-}\Bigg)Y_{lm}+\dots$
$=\Bigg[\sum_{l,m}\phi_{R(l,m)}^{-}\!\left(\sum_{k=0}^{l}\frac{(-1)^{k}}{2^{k}(k!)^{2}}\frac{(l+k)!}{(l-k)!}\right)Y_{lm}\Bigg]\frac{\log r}{r}$
$\quad+\Bigg[\sum_{l,m,k\leq l}\frac{(-1)^{k}}{2^{k}(k!)^{2}}\frac{(l+k)!}{(l-k)!}\big(\varphi_{R(l,m)}^{-}-\phi_{R(l,m)}^{-}H_{k}\big)Y_{lm}\Bigg]\frac{1}{r}+\dots\,.$
The above expressions can be simplified using the identity (93)
$\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}=P_{l}(0).$
The complete expression near $i^{0}$ is thus the sum of the advanced and retarded terms, which gives
$\left.\phi\right|_{i^{0}}$ $=\left[\sum_{l,m}P_{l}(0)\left(\phi_{A\left(l,m\right)}^{+}+\phi_{R\left(l,m\right)}^{-}\right)Y_{lm}\right]\frac{\log r}{r}+$
$+\left[\sum_{l,m}\Bigg(P_{l}(0)\left(\varphi_{A\left(l,m\right)}+\varphi_{R\left(l,m\right)}^{-}\right)\right$. (70)
$\hskip 28.45274pt-\left.\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}H_{k}\left(\phi_{A\left(l,m\right)}^{+}+\phi_{R\left(l,m\right)}^{-}\right)\Bigg)Y_{lm}\right]\frac{1}{r}+\dots\,.$
Let us now compute the canonical momentum $\pi=\partial_{t}\phi$ near $i^{0}$. The contribution from the advanced solution is given by
$\left.\pi_{A}\right|_{i^{0}}$ $=\left[\sum_{l,m}\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\right)\phi_{A\left(l,m\right)}^{+}Y_{lm}\right]\frac{\log r}{r^{2}}$
$+\left[\sum_{l,m}\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\phi_{A\left(l,m\right)}^{+}\left(1-kH_{k}\right)+k\varphi_{A\left(l,m\right)}\right)\right)Y_{lm}\right]\frac{1}{r^{2}}+\dots\,.$
The retarded component of the momentum near $i^{0}$ takes the form
$\left.\pi_{R}\right|_{i^{0}}=$ $-\left[\sum_{l,m}\sum_{k=1}^{l}\frac{\left(-1\right)^{k}k}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\phi_{R\left(l,m\right)}^{-}Y_{lm}\right]\frac{\log r}{r^{2}}+$
$+\left[\sum_{l,m}\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\phi_{R\left(l,m\right)}^{-}\left(1+kH_{k}\right)-k\varphi_{R\left(l,m\right)}^{-}\right)Y_{lm}\right]\frac{1}{r^{2}}+\dots\,.$
This expression can be simplified by applying the following identity (94)
$\sum_{k=0}^{l}\frac{\left(-1\right)^{k}k}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}=-P_{l}^{\prime}(0),$
Thus, the combined contribution of the retarded and advanced components of the momentum near $i^{0}$ becomes
$\left.\pi\right|_{i^{0}}$ $=-\left[\sum_{l,m}P_{l}^{\prime}(0)\left(\phi_{A\left(l,m\right)}^{+}-\phi_{R\left(l,m\right)}^{-}\right)Y_{lm}\right]\frac{\log r}{r^{2}}$
[…]
$+\left.\left.\left(\sum_{k=0}^{l}\frac{\left(-1\right)^{k}}{2^{k}\left(k!\right)^{2}}\frac{\left(l+k\right)!}{\left(l-k\right)!}kH_{k}\right)\left(\phi_{R\left(l,m\right)}^{-}-\phi_{A\left(l,m\right)}^{+}\right)\right)Y_{lm}\right]\frac{1}{r^{2}}+\dots\,$
As noted in Fuentealba:2024lll , terms of order $\log r/r$ in $\left.\phi\right|_{i^{0}}$ and $\log r/r^{2}$ in $\left.\pi\right|_{i^{0}}$ are not allowed. Therefore, from Eqs. (70) and (71), we obtain the following conditions
$\sum_{l,m}P_{l}(0)\left(\phi_{A\left(l,m\right)}^{+}+\phi_{R\left(l,m\right)}^{-}\right)Y_{lm}$ $=0\,$, (72)
$\sum_{l,m}P^{\prime}_{l}(0)\left(\phi_{A\left(l,m\right)}^{+}-\phi_{R\left(l,m\right)}^{-}\right)Y_{lm}$ $=0\,.$ (73)
There are two cases depending if $l$ is odd or even. Thanks to the parity properties of the Legendre polynomials, we have that half of the equations are always satisfied. So, for $l$ even, only the first equation imposes a condition on the field coefficients. In contrast, for $l$ odd only the second relation yields a condition on the difference of the coefficients $\phi^{\pm}_{(l,m)}$.
Consequently, the condition of not having logarithmic terms in (70) and (71) imposes the following relations between the coefficients $\phi_{A\left(l,m\right)}^{+}$ and $\phi_{R\left(l,m\right)}^{-}$ that can be summarized as
$\phi_{A\left(l,m\right)}^{+}=-\left(-1\right)^{l}\phi_{R\left(l,m\right)}^{-}\,.$ (74)
The above expression establishes a relation between the leading logarithmic contributions of the field at $\mathscr{I}_{-}^{+}$ and $\mathscr{I}_{+}^{-}$. Using the definitions of the radial logarithmic terms given in Eqs.(68) and (69), and imposing the condition in (74), one obtains:
[…]
$=0\,.$
[…]
which corresponds precisely to the matching condition established in Fuentealba:2024lll .
If we now assume that there is no leading logarithmic term ($\Psi(\hat{x})=\Psi^{\prime}(\hat{x})=0$), the solution reduces at leading order to the $P$-branch (in the limit $u\rightarrow$-$\infty$, $v\rightarrow\infty$), which obeys the “standard” parity condition of Campiglia:2017dpg .
Appendix B General solution of Maxwell’s equations in retarded coordinates
The wave equation obeyed by the electromagnetic potential is
$\left(\boxempty A_{\mu}-\partial_{\mu}\nabla_{\nu}A^{\nu}\right)=0.$
Writing this solution in terms of retarded coordinates, and imposing for simplicity the gauge $A_{r}=0$, one finds the following equations:
$\partial_{r}^{2}A_{u}+\frac{2}{r}\partial_{r}A_{u}-\partial_{u}\partial_{r}A_{u}+\frac{1}{r^{2}}\left(D^{2}A_{u}-\partial_{u}\left(D^{B}A_{B}\right)\right)$ $=0\,$, (75)
$\partial_{r}^{2}A_{u}+\frac{2}{r}\partial_{r}A_{u}-\frac{1}{r^{2}}\partial_{r}\left(D^{B}A_{B}\right)$ $=0\,$, (76)
$\partial_{r}^{2}A_{A}-2\partial_{u}\partial_{r}A_{A}+\partial_{A}\partial_{r}A_{u}+\frac{D^{B}D_{B}A_{A}-D_{A}D_{C}A^{C}-A_{A}}{r^{2}}$ $=0\,.$ (77)
To solve these equations, it is convenient to introduce the following separation of variables
$A_{u}$ $=f_{lm}\left(u,r\right)Y_{lm}\,$,
$A_{A}$ $=g_{lm}\left(u,r\right)\partial_{A}Y_{lm}+h_{lm}\left(u,r\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\,.$
Thus, from Eq. (75) one finds
$\partial_{r}^{2}f_{lm}+\frac{2}{r}\partial_{r}f_{lm}-\partial_{u}\partial_{r}f_{lm}+\frac{1}{r^{2}}l\left(l+1\right)\left(\partial_{u}g_{lm}-f_{lm}\right)=0.$
From Eq. (76) one obtains
$\partial_{r}^{2}f_{lm}+\frac{2}{r}\partial_{r}f_{lm}+\frac{1}{r^{2}}l\left(l+1\right)\partial_{r}g_{lm}=0\,.$
From Eq. (77) one finds
$\left(\partial_{r}^{2}g_{lm}+\partial_{r}f_{lm}-2\partial_{u}\partial_{r}g_{lm}+\frac{1}{r^{2}}l\left(l+1\right)g_{lm}-\frac{1}{r^{2}}g_{lm}\right)\partial_{A}Y_{lm}\\ +\left(\partial_{r}^{2}h_{lm}-2\partial_{u}\partial_{r}h_{lm}-\frac{1}{r^{2}}h_{lm}\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\\ +\frac{1}{r^{2}}g_{lm}D^{C}D_{A}D_{C}Y_{lm}+\frac{1}{r^{2}}h_{lm}\sqrt{\gamma}\epsilon_{AB}D^{C}D^{B}D_{C}Y_{lm}=0.$
Using the identity $D^{B}D_{A}D_{B}\Lambda=D_{A}D^{2}\Lambda+D_{A}\Lambda$, one obtains
$\left(\partial_{r}^{2}g_{lm}+\partial_{r}f_{lm}-2\partial_{u}\partial_{r}g_{lm}\right)\partial_{A}Y_{lm}+\\ \left(\partial_{r}^{2}h_{lm}-2\partial_{u}\partial_{r}h_{lm}-\frac{l\left(l+1\right)}{r^{2}}h_{lm}\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}=0\,.$ (78)
This gives two equations
$0=\partial_{r}^{2}g_{lm}+\partial_{r}f_{lm}-2\partial_{u}\partial_{r}g_{lm}\,$,
$0=\partial_{r}^{2}h_{lm}-2\partial_{u}\partial_{r}h_{lm}-\frac{l\left(l+1\right)}{r^{2}}h_{lm}\,.$
In summary, Maxwell equations reduce to the following system of differential equations
Equations for $f$ and $g$:
$0$ $=\partial_{r}^{2}f_{lm}+\frac{2}{r}\partial_{r}f_{lm}-\partial_{u}\partial_{r}f_{lm}+\frac{1}{r^{2}}l\left(l+1\right)\left(\partial_{u}g_{lm}-f_{lm}\right)\,$, (79)
$0$ $=\partial_{r}^{2}f_{lm}+\frac{2}{r}\partial_{r}f_{lm}+\frac{1}{r^{2}}l\left(l+1\right)\partial_{r}g_{lm}\,$, (80)
$0$ $=\partial_{r}^{2}g_{lm}+\partial_{r}f_{lm}-2\partial_{u}\partial_{r}g_{lm}\,.$ (81)
Equation for $h$:
$0=\partial_{r}^{2}h_{lm}-2\partial_{u}\partial_{r}h_{lm}-\frac{l\left(l+1\right)}{r^{2}}h_{lm}\,.$ (82)
We shall assume, as in the scalar case, that the functions $f_{lm}$, $g_{lm}$ and $h_{lm}$, can be decomposed as the sum of retarded and advanced waves. The retarded part is assumed to have an asymptotic expansion in inverse powers of $r$ with coefficients that are smooth functions of $u$ and of the angles. Similarly, the advanced part is assumed to have an asymptotic expansion in inverse powers of $r$ with coefficients that are smooth functions of $v$ and of the angles.
B.1 Solution of the magnetic sector
By performing the change of variables $h_{lm}\left(u,r\right)=r\phi\left(u,r\right)$ in Eq. (82), the differential equation reduces to that of a scalar field (Eq. (5)). Therefore, the general solution to Eq. (82) is given by
$h_{lm}\left(u,r\right)=\sum_{k=0}^{l}\frac{1}{k!\left(2r\right)^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\frac{d^{l-k}}{du^{l-k}}h_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}h_{lm}^{\text{adv}}\left(v\right)\right)\,.$ (83)
B.2 Solution of the electric sector
The equations that must be solved are given by the following system of differential equations:
$0$ $=\partial_{r}^{2}f_{lm}+\frac{2}{r}\partial_{r}f_{lm}-\partial_{u}\partial_{r}f_{lm}+\frac{1}{r^{2}}l\left(l+1\right)\left(\partial_{u}g_{lm}-f_{lm}\right)\,$, (84)
$0$ $=\partial_{r}^{2}f_{lm}+\frac{2}{r}\partial_{r}f_{lm}+\frac{1}{r^{2}}l\left(l+1\right)\partial_{r}g_{lm}\,$, (85)
$0$ $=\partial_{r}^{2}g_{lm}+\partial_{r}f_{lm}-2\partial_{u}\partial_{r}g_{lm}\,.$ (86)
If we multiply (85) by $r^{2}$ we find an exact differential that integrates as follows:
$r^{2}\partial_{r}f_{lm}+l\left(l+1\right)g_{lm}=\alpha_{lm}\left(u\right)l\left(l+1\right).$ (87)
Replacing (87) in (86) we find the following equation for $g_{lm}$:
$\partial_{r}^{2}g_{lm}-2\partial_{u}\partial_{r}g_{lm}-\frac{l\left(l+1\right)}{r^{2}}g_{lm}=-\frac{l\left(l+1\right)\alpha_{lm}\left(u\right)}{r^{2}}.$
Note that this equation is identical to that of the magnetic sector in (82) but with a source given by $\alpha_{lm}\left(u\right)$. Therefore, the homogeneous solution will be identical to the one of the magnetic sector in (83). On the other hand, a particular solution is given by $g_{lm}^{\text{part}}=\alpha_{lm}$. Consequently, the general solution is given by
$g_{lm}\left(u,r\right)=\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)+\alpha_{lm}\left(u\right)\,$, (88)
for $l\geq 1$. Note that the zero mode is irrelevant because it enters in the term with $\partial_{A}Y_{lm}$. For future convenience, we have added a minus sign in front of $g_{lm}^{R}\left(u\right)$.
In order to determine $f_{lm}\left(u,r\right)$, let us substract equations (84) and (85)
$0=-\partial_{u}\partial_{r}f_{lm}+\frac{1}{r^{2}}l\left(l+1\right)\left[\partial_{u}g_{lm}-f_{lm}-\partial_{r}g_{lm}\right].$
Taking a $u$-derivative of Eq. (87) and replacing it in the previous equation, we find
$0=\frac{1}{r^{2}}l\left(l+1\right)\left[2\partial_{u}g_{lm}-\partial_{r}g_{lm}-\partial_{u}\alpha_{lm}-f_{lm}\right].$
Note that for $l\geq 1$, one obtains the following expression for $f_{lm}$
$f_{lm}\left(u,r\right)=2\partial_{u}g_{lm}\left(u,r\right)-\partial_{r}g_{lm}\left(u,r\right)-\partial_{u}\alpha_{lm}\left(u\right)\qquad\qquad\text{for }l\geq 1.$
The case $l=0$ must be treated separately. Indeed, for $l=0$, the equations (84)-(86) reduce to
$0$ $=\partial_{r}^{2}f_{00}+\frac{2}{r}\partial_{r}f_{00}-\partial_{u}\partial_{r}f_{00}\,$, (89)
$0$ $=\partial_{r}^{2}f_{00}+\frac{2}{r}\partial_{r}f_{00}\,$, (90)
$0$ $=\partial_{r}^{2}g_{00}+\partial_{r}f_{00}-2\partial_{u}\partial_{r}g_{00}\,.$ (91)
In this case, the equations for $f_{lm}$ decouple automatically. From Eqs. (89) and (90) we find
$f_{00}=\frac{Q}{r}-\partial_{u}\alpha_{00}\left(u\right)\,.$
The zero mode $g_{00}$ can be determined from Eq. (91), but is completely irrelevant for the analysis because it appears in the term with $\partial_{A}Y_{lm}$.
Consequently, the expression for $f_{lm}\left(u,r\right)$ is the following:
$f_{lm}\left(u,r\right)$ $=-\sum_{k=0}^{l}\frac{1}{2^{k-1}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\frac{d^{l-k+1}}{du^{l-k+1}}g_{lm}^{\text{ret}}\left(u\right)$
$+\sum_{k=0}^{l}\frac{k}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)+\partial_{u}\alpha_{lm}\left(u\right)\,.$
The expression for the electromagnetic potential is then given by
$A_{u}$ $=-\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k-1}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\frac{d^{l-k+1}}{du^{l-k+1}}g_{lm}^{\text{ret}}\left(u\right)Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{k}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)Y_{lm}+\frac{Q}{r}+\partial_{u}\alpha\,$,
$A_{A}$ $=\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)\partial_{A}Y_{lm}+\partial_{A}\alpha$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\frac{d^{l-k}}{du^{l-k}}h_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}h_{lm}^{\text{adv}}\left(v\right)\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm},$
where $\alpha$ is function of $u$ and $\hat{x}$.
It is convenient to fix the gauge such that near $\mathscr{I}^{+}$ one has $A_{u}=O\left(r^{-1}\right)$. This can be done by choosing
[…]
and is compatible with $A_{r}=0$ since $\alpha$ does not depend on $r$.
With this gauge choice, we finally get for the full components (retarded + advanced) of the electromagnetic potential in retarded coordinates the following expressions,
$A_{u}$ $=-\sum_{l,m}\sum_{k=1}^{l}\frac{1}{2^{k-1}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\frac{d^{l-k+1}}{du^{l-k+1}}g_{lm}^{\text{ret}}\left(u\right)Y_{lm}$
$+\sum_{l,m}\sum_{k=0}^{l}\frac{k}{2^{k}k!r^{k+1}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)Y_{lm}+\frac{Q}{r}\,.$
$A_{A}$ $=\sum_{l,m}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(-\frac{d^{l-k}}{du^{l-k}}g_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}g_{lm}^{\text{adv}}\left(v\right)\right)\partial_{A}Y_{lm}$
$+2\sum_{l,m}\frac{d^{l}}{du^{l}}g_{lm}^{\text{ret}}\left(u\right)\partial_{A}Y_{lm}$
$+\sum_{lm}\sum_{k=0}^{l}\frac{1}{2^{k}k!r^{k}}\frac{\left(l+k\right)!}{\left(l-k\right)!}\left(\frac{d^{l-k}}{du^{l-k}}h_{lm}^{\text{ret}}\left(u\right)+\left(-1\right)^{k}\frac{d^{l-k}}{dv^{l-k}}h_{lm}^{\text{adv}}\left(v\right)\right)\sqrt{\gamma}\epsilon_{AB}\partial^{B}Y_{lm}\,.$
Appendix C Legendre polynomials properties
In this article, multiple sums appeared, some of which have known results. In this section, we will show how to compute them.
There are two types of sums: those appearing at null infinity, which yield the matching conditions, and those arising when evaluating the field at spatial infinity. Both can be expressed in terms of Legendre polynomials and their derivatives, evaluated at specific points, either x=1x=-1 or $x=0$.
C.1 Values at $x=0$
An interesting representation of the Legendre polynomials is given by
$P_{l}(x)=\sum_{k=0}^{l}\frac{(x-1)^{k}}{2^{k}(k!)^{2}}\frac{(l+k)!}{(l-k)!};\quad\text{with}\quad P_{l}(x)=(-1)^{l}P_{l}(-x).$ (92)
This representation is useful because several of the relevant sums can be written as Legendre polynomials and their derivatives evaluated at zero:
$P_{l}(0)=$ $\sum_{k=0}^{l}\frac{(-1)^{k}}{2^{k}(k!)^{2}}\frac{(l+k)!}{(l-k)!};\quad P_{l}(0)=0\text{ for }l\text{ odd}$ (93)
$-P^{\prime}_{l}(0)=$ $\sum_{k=0}^{l}\frac{k(-1)^{k}}{2^{k}(k!)^{2}}\frac{(l+k)!}{(l-k)!}\quad P^{\prime}_{l}(0)=0\text{ for }l\text{ even}$ (94)
$P^{\prime\prime}_{l}(0)=$ $\sum_{k=0}^{l}\frac{k(k-1)(-1)^{k}}{2^{k}(k!)^{2}}\frac{(l+k)!}{(l-k)!}\quad P^{\prime\prime}_{l}(0)=0\text{ for }l\text{ odd}.$ (95)
Each of these sums arises when evaluating the fields at $i^{0}$. The appearance of these polynomials is not surprising, as they naturally occur in Fuentealba:2024lll ; Fuentealba:2025ekj .
C.2 Values at $x=\pm 1$
The other type of sums in this article are those related to the field evaluated at null infinity (either future or past). In general, they can always be expressed as Legendre polynomials evaluated at x=1x=-1. From the same representation in (92), we can write
$P_{l}(-1)=$ $\sum_{k=0}^{l}\frac{(-1)^{k}}{(k!)^{2}}\frac{(l+k)!}{(l-k)!}$ (96)
$-2P^{\prime}_{l}(-1)=$ $\sum_{k=0}^{l}\frac{k(-1)^{k}}{(k!)^{2}}\frac{(l+k)!}{(l-k)!}.$ (97)
Due to the parity condition and the representation of the Legendre polynomials in (92) we can easily determine the value of $P_{l}(-1)$ and their derivatives
$P_{l}(-1)=(-1)^{l}P_{l}(1)=(-1)^{l}\,,\quad-2P^{\prime}_{l}(-1)=2(-1)^{l}P^{\prime}_{l}(1)=(-1)^{l}l(l+1).$ (98)